### v1

# Creative and geometric times in physics, mathematics, logic, and philosophy

Abstract
We propose a distinction between two different concepts of time that play a role in physics: geometric time and creative time. The former is the time of deterministic physics and merely parametrizes a given evolution. The latter is instead characterized by real change, i.e. novel information that gets created when a non-necessary event becomes determined in a fundamentally indeterministic physics. This allows us to give a naturalistic characterization of the present as the moment that separates the potential future from the determined past. We discuss how these two concepts find natural applications in classical and intuitionistic mathematics, respectively, and in classical and multivalued tensed logic, as well as how they relate to the well-known A- and B-theories in the philosophy of time.
I Introduction
There is almost nothing that we perceive so ubiquitously than the passage of time. And yet our most successful physical theories still struggle to make sense of this concept in an unequivocal way. Actually, modern physics has relegated time to play a less and less special role [1]. However, in the words of I. Prigogine, ``````` `no formulation of the laws of nature that does not take into account this constructive role of time can ever be satisfactory” [2].
To address this fundamental problem, we propose here two different concepts of time, which we call geometric time and creative time, respectively. We show that those stem from our fundamental assumption of physics as being either deterministic or indeterministic at the fundamental level, and that they both seem to contribute to our understanding of natural phenomena. In particular, we will see that geometric time is the parametric time appearing in the (deterministic) equations of motions of physics, just a coordinate that, together with the three spatial directions, labels fix events in a geometrical block-universe. On the other hand, creative time originates from the assumption that certain events which are fundamentally indeterminate become determinate. Our general approach is admittedly highly influenced by quantum mechanics, which has brought the concept of fundamental indeterminacy into the domain of physics. But we abstractly conceive the possibility that there is ontic indeterminacy in the world (also at the classical level, see [3, 4]). This idea of creative time resonates to some extent with the positions of the philosopher H. Bergson, well-known for having confronted Einstein about the nature of time, who stated: ``````` `Time is what prevents everything from being given all at once. […] It must therefore be development. Wouldn’t it then be the vehicle of creation and choice? Would time’s existence not prove that things are undetermined? Would time not be this very indeterminacy?” [5]. We somewhat twist this view around by assuming that it is indeterminacy that defines a different fundamental type of time. Indeed, if there is ontic indeterminacy in nature (i.e., physical pure states do not fully determine the future), this leads to multiple potential future states, i.e., to indeterminism and an open future. It is the actualization of these potentialities, i.e., when new information previously not existing is created, that makes this creative time tick.
Figure 1: Typical evolution of a classical chaotic system in phase space. The initial state (a) has some ontic indeterminacy, i.e., the position and momentum are in this case not determined further than the finite region depicted in figure. When evolved through the deterministic equations of motions (b), the indeterminacy spreads (while preserving the same volume), leading to multiple potential future evolutions, i.e. to indeterminism. However, only one of these potential evolutions will actualize (c), while ruling out the other previously potential futures (dashed); it is this process of actualization that makes creative time process.
Starting from physics (Sect. II), we apply these two concepts of time to different disciplines, showing that this distinction gives rise to natural parallels in mathematics (classical versus intuitionistic; Sect. III), helps clarify the logic of scientific propositions (Sect. IV), and it contributes further to the metaphysics of time (so-called A-theory versus B-theory; Sect. V). More specifically, the approach based on consecutive actualizations of potentialities, that is when creative time is at work, leads to a naturalistic characterization of the difference between past, present and future.
II Two times in physics
In order to illustrate different concepts of time in physics, let us start from an example. Let us ask the following scientific question: what will be the weather like, say on the field hockey pitch of Geneva, exactly in one month from now? One can envision several procedures to reply to this question. On the one hand, it is possible to use the equations of motion of all the involved molecules of air, water vapour, etc., to calculate their evolution and so predict the weather in Geneva in a month time. On the other hand, one can wait exactly one month, let the process develop, and finally observe the weather in Geneva. In-between, one could spend one week to collect more data – or merely waste a week – and only then start computing the evolution of all air molecules. Let us concentrate on the two first mentioned extremal procedures. The question that we ask here is: Are these two procedures talking about the same unique concept of time, or are there multiple ones that play a role in physics? We contend that there are (at least) two different concepts of time, as exemplified by the two procedures above.
Before elaborating on that, we ought to distinguish between different fundamental interpretations of physics, which become especially relevant in dealing with complex systems like the one determining the weather in our example. The standard narrative (i) has it that classical physics is deterministic, i.e. that each state is mapped one-to-one to all other possible states at all times (with time being the parameter that appears in the equations of motion). Although usually not stressed enough, this is based on two independent ingredients: the first is that since the equations of motion are ordinary differential equations, their solutions generally have a unique continuation in the parameter time, [^1] In exceptional cases, classical equations of motion of even simple systems do not have a unique (deterministic) solution, such as in the example of Norton’s dome [6]. However, in highly complex systems – such as Lagrangian fluid particles in high Reynolds-number turbulence – the fact that deterministic trajectories are not unique is rather common (see [7] and references therein). given appropriate initial conditions; the second is that the initial conditions exist fully determinate with infinite precision (i.e., a state is represented by a dimensionless mathematical point in phase space, in turn described by an n-tuple of real numbers). Only taking these two assumptions together justifies the purported fundamental determinism of classical physics. Another possibility, (ii) is to keep the idea that the system is fundamentally deterministic, but realistically it is not possible to know the initial conditions with infinite precision. This becomes most relevant in chaotic systems (like the weather in our example) in which the future behavior of the considered system highly depends on the infinitesimal digits of the initial conditions. In this case one can only statistically predict what are the possible future states compatible with the partial knowledge of the initial conditions. Finally, (iii) one can instead uphold the more radical assumption that the initial conditions (alternatively, the state of a system) simply do not exist determined with infinite precision, but there is an ontic indeterminacy in nature (for a collection of works on this see, [8, 3, 9, 10, 11, 4]; see also [12] and references therein for recent developments on metaphysical indeterminacy). In this case, although the equations of motion may remain unchanged, the system does not undergo a necessary evolution, but multiple future states are possible (perhaps with a determined objective tendency or propensity; see [4]), i.e., it is the ontic indeterminacy that leads to fundamental indeterminism (see Fig. 1). In what follows, when referring to indeterminacy we will have this kind of ontic property in mind and its consequent indeterminism.
Let us now go back to analyze the concepts of time that are at work in the example of the weather in Geneva in a month. The first procedure, under the ideal assumption of determinism (i), merely makes use of time as a parameter that simply indicates the coordinate of a certain evolution (e.g., sunny weather in one month from the start of the computation). Together with the three spatial coordinates (i.e., in Geneva in our example), this kind of time forms a manifold that contains all the events: a four-dimensional block universe picture where past, present and future are all on the same metaphysical footing (this is also known as Eternalism in the philosophy of time; see Sect. V). This can thus be called geometric time, the time that most physicists have come to think of as fundamental, both from classical physics and even more from relativity theory where time becomes “just” another dimension like space. In this sense, time is a relational label that allows one to position an event (with respect to another point of reference), but there is no real ``````` `becoming”.
On the other hand, let us consider the case of ontic indeterminacy in the physical state (iii). There, the actual development of a non-necessary physical process – such as the evolution of a highly chaotic system like the one that determines the weather, or of a quantum system – involves, in our opinion, a time that really ``````` `processes”. That is, a concept of time which sets an actual separation of the future from the past (with the present at their interface). And it does so when new information that was not existing before gets created – i.e., even having access to the whole existing information of the initial conditions (even of the whole state of the universe) before, would not determine all the future physical properties of a system afterwards. Again, we stress that we are not talking here about epistemic uncertainty, but rather we explicitly take a physicalist stance on information: it quantifies something objectively encoded in the degrees of freedom of the universe. In this sense information gets created when non-necessary physical events come about, independently of any observer. Note that, – and this is what happens also in quantum physics – although the state of the system is fully determined and it contains the maximal information about the systems (i.e., it is pure), this does not fully determine the future evolution (see also Ref. [13]). The concept of time that stems from this can thus be called creative time. [^2] The distinction between geometric and creative time was sketched in Refs. [14, 15, 3, 16] and previously also in Ref. [17] where the creative time was called ``````` `historical time”. On a similar note, G. Ellis remarked that ``````` `The most important property of time is that it unfolds. […] The time that is the present at this instant will be the past at the next instant. […] It is this fundamental feature of time that is not encapsulated by today’s theoretical physics” [18] (see also [19]).
Let us emphasize that the concept of creative time has nothing to do with language, nor with subjective or psychological processes; it ticks due to objective processes that happen by themselves in nature. Accordingly, there are two types of events: necessary events (e.g., when two geometric worldlines cross each other) and non-necessary events when potentialities actualize. Those instances of actualization determine the objective present. In other words, this fundamental “becoming” comes from the realization of one single outcome among the possible ones (with the contextual exclusion of the other previously valid futures) due to ontic indeterminacy.
In the literature one often finds that time ``````` `unfolds” or that time ``````` `flows” to indicate that time is associated with something that really changes (so with creative time). Here we prefer to use the terms time ``````` `develops” or ``````` `processes”, because also a film already shot does unfold, but nothing new is created there; whereas to flow is a somewhat misleading term for it reminds of the concept of speed which has time already in it. So, henceforth, we will write that geometric time passes while creative time processes or happens.
Coming back to the example of the weather in Geneva in a month, to carry out either of the envisioned procedures to determine the outcome takes some time. In the first procedure, i.e. when geometric time is at work, one maps the time that it takes to perform the computation $\Delta t^{\prime}$ to the time appearing in the equations of motion. That is, to the interval $\Delta t=t_{f}-t_{i}$, where $t_{i}$ is the parametric time associated to the initial state and $t_{f}$ is the parametric time of the considered final state (i.e., the moment when we want to know the weather in Geneva). The (deterministic) computation lasts $\Delta t^{\prime}<\Delta t=1$ month, and returns as an outcome the weather in Geneva at $t_{f}$, say ``````` `sunny”. [^3] In principle the time for the “prediction” could take longer than the time interval to observe the outcome. In such a case, one would anyways ask for consistency of the calculation afterwards.
One can also think of more subtle questions about the weather, such as ``````` `will there be (anytime in history) 700 consecutive days of rain in Geneva?”. If the complex system that forms the weather is fully deterministic, answering this question seems only a matter of geometric time, although the answer may remain undecided for any finite value of geometric time. One can run the computation of the dynamics of the system and at some point find that the answer is ``````` `yes”, if one finds the 700 consecutive days of rain, otherwise the algorithm will keep searching without ever halting (however, this question might be more complex, see Sect. III).
Two remarks are here in order. Firstly, note that $\Delta t^{\prime}$ can be speed up by a more efficient computation, namely, this depends on how fast one ``````` `runs” through the parameter time as it appears in the equations of motion (this depends on the used algorithm and, physically, one can really think of the speed of the head of a Turing machine in reading the tape). Moreover, time dilation in relativity can be seen as a physical process to modify the interval of geometric time (see Subsect. VI.1). So the map $\Delta t\rightarrow\Delta t^{\prime}$ is a one-to-many function. Secondly, once one reaches the final state with the associated outcome ``````` `sunny”, one gains full information about the initial state at $t_{i}$ and its consequences, but not necessarily the full information about the final state if there is some fundamental indeterminacy, i.e., the actual weather that will be observed may vary.
On the contrary, the procedure involving creative time, namely, of waiting a month and carrying out an observation of the weather gives full information about the outcome, but provides in general little information about the initial state. This is because new information has been created along the way, through a series of indeterministic processes. So, the initial and the final states are not deterministically mapped to each other. The exactly same initial state under exactly the same physical conditions leads in general to different outcomes, so one can only infer a set of initial states that are compatible with the observed outcome. Creative time is thus associated to the possibility that physical properties are in general ontologically indeterminate (e.g., it is neither true that ``````` `the particle is within region $r$” nor that ``````` `the particle is not within $r$”, where $r$ is a region of space, see also Sect. IV). Creative time is required when an actualization event of the values of physical variables that were previously indeterminate occurs. It is the process of change from indeterminate to determinate – with an associated creation of novel information – that brings about this “happening” of time: creative time ticks when new information is created.
III Two times in Mathematics
Although it is not common to discuss time in mathematics at all, the problem of the two different kinds of time introduced above has almost a perfect counterpart in mathematics. Let us now consider a standard mathematical object, e.g. a real number. Typical real numbers – in fact, almost all of them in a mathematical sense – are uncomputable, i.e. for each of these numbers there exists no algorithm – which is by definition a finite list of instructions – that outputs all their digits (they have infinite Kolmogorov complexity) [10]. On the contrary, all the (irrational) numbers that we usually consider as prototypical examples of reals, such as $\sqrt{2}$ or $\pi$, are all computable (the ratio between the diagonal of a square to its edge is a simple algorithm that outputs all the digits of $\sqrt{2}$ and similarly the ratio of the circumference to the diameter of a circle for $\pi$). Obviously, all the rational numbers are also computable: they can be directly written down or compressed (in the case of repeating decimals) in a finite string.
The standard way of regarding mathematics, the so-called classical-mathematics, is a form of Platonism which posits that, among the other mathematical entities, real numbers exist with their infinite series of digits, although there is in general no way to even label and thus grasp them (since there is no algorithm that generates them [20, 21]. Hence, classical Platonistic mathematics is a timeless language, which per se is not a bug, but it becomes problematic when mathematics is elevated to the language of science, i.e., it is used to the describe the physical world.
However, there are alternative approaches known as constructive mathematics, of which the most prominent is intuitionism. Therein, mathematical entities are not given at once, but rather are processes in development (digits get created continuously, one after the other). This provides mathematics with a concept of passage of time (and it would be only at infinite time – at the end of time, so to say – that the mathematical entities, such as real numbers, are completed into the ones defined by classical mathematics). The initiator of intuitionism, L.E.J. Brouwer, envisioned a ``````` `creative subject”, i.e., an idealized mind or mathematician, who is responsible for this progressive process of creation in time of, for instance, the digits of an uncomputable real numbers. Several authors, however, have distanced themselves from this controversial concept of a creative subject. In particular, one of us (N.G.) has put forward the idea that the digits of a typical real number are generated by a true random number generator, [^4] More precisely, it is not necessarily new digits that come into existence, but the new information reduces the indeterminacy of intuitionistic real numbers [21]. i.e., a natural process that is able to create a piece of new information by changing a fundamentally indeterminate bit into a determined one [10]. [^5] For more works that relate constructive mathematics to physics, see [20] and [22]. This version of intuitionism could thus be labeled ``````` `objective” or ``````` `naturalistic” intuitionism. In this way, typical real numbers become a graspable concept directly linked to creative time, that is, to the change from the indeterminate to the determinate. Hence, naturalistic intuitionism, is a tensed mathematical language.
To think about computable real numbers, one just needs geometric time. In fact, it is possible to think that the full information about those numbers is contained in the (finite) algorithm that defines them, i.e., in their initial conditions. For instance, one can ask what is the 43800th digit of $\pi$ and the answer is given by running an algorithm that outputs the digits of $\pi$ and picking its 43800th digit. Note that it is not necessary to go through all the previous digits of $\pi$ to compute 43800th one. There are so-called digit-extraction algorithms that allow one to directly compute the nth digit of $\pi$ [23]. [^6] The most well-known of such algorithms is perhaps the Bailey-Borwein-Plouffe formula [24], that allows to directly compute the hexadecimal digits of $\pi$. Digit-extraction algorithms, also in base 10, are known for several other irrational computable real numbers, such as for $e$. This can be seen as a further evidence that computable (irrational) numbers are already fully determined and do not require creative time. The point is that since this is found through a deterministic outcome, the answer to this kind of questions is fully contained in the algorithm, which can be run more or less fast, giving the ability to manipulate this ``````` `mathematical geometric time”. This exactly resembles the weather example in Sect. II for the deterministic evolution of the physical system that forms the weather.
In contrast, if one considers instead a typical, i.e., uncomputable, real number, things are different. There are no algorithms compressing the information of that number. The only existing ``````` `algorithm” is the number itself. Each next digit is generated by a genuinely natural random process, therefore, asking which is the 43800th digit of such a number requires to wait for all 43800 instances of creating of the ``````` `next digit” (see footnote 4), i.e., it requires ``````` `mathematical creative time”.
Finally, one can also ask a question of the kind ``````` `are there 700 consecutive sevens in the digits of a number. [^7] The example of 700 consecutive sevens is borrowed from C. Posy [25]. If that number is uncomputable, then one has no choice but to wait and see. However, even for computable numbers, like e.g. $\pi$, the question is interesting. The situation is similar to the example of the weather – i.e., 700 consecutive days of rain if one assumes determinism – but the axiomatic character of mathematics makes the solution to this question clearer. One can, in fact, think of programming two softwares: The first, $S_{1}$ is exactly the one we used in the physical example of the weather, i.e., the systematic search of a 700 consecutive sevens in $\pi$; on the other hand, the second software, $S_{2}$ outputs all the theorems that can be derived within Peano arithmetic and it halts should it find the negation of the conjecture. If the statement is true, $S_{1}$ will find this sequence of 700 sevens and therefore halt after a finite (geometric) time. If the statement is provable (within Peano’s arithmetic), $S_{2}$ will halt after a finite (though presumably enormous) geometric time. Interestingly, if one assumes that neither halts, then the statement is false (if not $S1$ would halt), but not provable within Peano’s arithmetic (if not $S_{2}$ would halt). [^8] For the sake of completeness, there is also the fourth logical possibility that both $S_{1}$ and $S_{2}$ halt, a case which would prove Peano’s arithmetic inconsistent. This would be an example of Gödel’s celebrated theorem. Note that mathematicians don’t use softwares like $S_{1}$ and $S_{2}$, but use their creativity to find shortcuts to analyse such statements, i.e., de facto they use creative time. In the case of the weather in Geneva, if one assumes that the system is indeterministic, like uncomputable numbers, then there is no choice but wait and see. However, if one trusts the equations describing the weather evolution and the fully determined initial conditions, then one could run analogues of $S_{1}$ and $S_{2}$, as for computable numbers. Note however, that this assumes one can compress the initial conditions into a finite algorithm (hence into a computer which is necessarily finite).
Let us acknowledge that this distinction between geometric and creative time in mathematics distances our view from the most accepted position in intuitionism, such as that of C. Posy [25] or M. van Atten [26, 27], who consider all numbers, whether computable or not, as generated by new information that comes about, so that they all require creative time in our parlance.
IV Logic
[…]
Figure 2: Representation of ontic indeterminacy in phase space.
Now, the principle of non-contradiction imposes already that propositions about physics must be referred to a time. Indeed, if one states simpliciter that ``````` `the system is inside $r$” and ``````` `the system is not inside $r$”, this is a violation of the principle of non-contradiction, since we do observe in nature the same system being sometimes inside $r$ and sometimes not. One has then to refer to a time to qualify the propositions. If one adds ``````` `at time $t$” to both propositions, then the principle of non-contradiction holds because –at any time $t$ – only one of these propositions would be true and the other not. On the contrary, if one labels one proposition with ``````` `at time $t_{1}$” and the other with ``````` `at time $t_{2}$”, with $t_{1}\neq t_{2}$, it is possible that they are both true. Hence, the principle of non-contradiction applies only to propositions at the same time. Note that this time is not necessarily creative, but it could just be the label appearing in the equations of motion, i.e., geometric time.
Coming instead to a scenario where there is ontic indeterminacy, we now show that the law of the excluded middle fails. Consider for instance a model of physics (classical, quantum, or even post-quantum) in which a (pure) state is represented by a finite volume, $R$, in phase-space (as opposed to a dimensionless mathematical point). That is, there is not merely a lack of knowledge about the system (as it is the case in, e.g., statistical physics), but this state contains the complete information existing at present about all the properties of the system (see also [3, 4]). Since, as we have already argued, this ontic indeterminacy generally leads to indeterminism in the future evolution of a system, [^10] Except for integrable classical dynamical systems. statements about future properties are in general neither true nor false because they are yet to be determined. And it is this process of determination that requires creative time. In general, in the case of ontic indeterminacy, one needs to introduce a multivalued tensed logic according to which a proposition is either true or false only at a time $\tau\geq\tau_{0}$ when it has become determined, but its truth value was ``````` `indeterminate” (as a third logical value) before $\tau_{0}$ (for multivalued logic see [28, 29]). As an example, consider a true random number generator that outputs a (single determined) bit of information, say $b=0$. That event defines the (creative) time instant $\tau_{0}$. After such an event that brings into existence this new bit of information, the truth value of the proposition ``````` `b=0” is true and the proposition ``````` `b=1” is false, but before $\tau_{0}$ both propositions had a truth value ``````` `indeterminate”. This represents a breakdown of the principle of the excluded middle. Note that such a failure of the excluded middle is found in all constructive mathematics – e.g. intuitionism [25]. However, despite these many similarities, we emphasize that our approach differs substantially from ``````` `orthodox” intuitionism. First of all, as already remarked in Section III, we reject the Brouwerian notion of ``````` `creative subject” in intuitionistic logic, advocating for a naturalistic version of intuitionism in which new information is created in the universe by natural processes. Moreover, contrarily to what we have advocated, we stress that orthodox intuitionistic logic is not three-valued (nor multivalued for that matter) – as proven by Gödel [30]– and that logical propositions are typically timeless.
Finally, consider a system whose present state is represented by a finite volume in phase space $R$ (see Fig. 2). This means that although the state is fully determined (but not as a collection of real numbers [9]), it does not fully determine the future. In this case, a statement like ``````` `the system is located in phase space within $r\subset R$” may lead to think that the principle of the excluded middle fails also for statements about the present. However, such a statement is not indeterminate, but actually false, for it is simply not the case that the system is located inside $r$ in phase space. The only presently true statement about the state of the system is ``````` `the system is located at $R$ in phase space” and nothing more. On the contrary, the statement ``````` `the system will be (at some future time) located in phase space at $r$” is currently indeterminate, because it is not impossible that the system will become more localized (in phase space). Once the evolution of the system develops, the statement will become either true or false. This process is what requires creative time. Hence, the failure of the excluded middle concerns only future statements, where the future is identified by a change of creative time.
V Relation to the philosophy of time
Before discussing in detail how our proposal of creative and geometric times relates to some positions in the philosophy of time, we would like to stress that, as most physicists, we are not taking a Humean position. On the contrary, we believe that laws of nature are real causal connections (they have an ontological status) and that they guide the evolution of things of the world at different times. Note that these laws may not connects states at different times by necessity (determinism), but only state a propensity [4]. In either case, this puts an emphasis on the role of time in physical laws.
In the vast literature devoted to the philosophy of time (see, e.g., [31, 32] and references therein), one finds a main separation into two camps, following the seminal works of J. M. E. McTaggart [33]. The two camps are customarily referred to as A-theory and B-theory of time. While we acknowledge that the distinction in A- and B-theories presents a variety of modification and it may not even encapsulate the most striking features of time in physics (e.g., directionality, openness of the future, objective becoming, etc.), we cannot avoid drawing some parallel with our distinction between creative and geometric time, respectively.
In fact, starting in reversed order, B-theorists regard positioning in time in dyadic relational terms: given two events $E_{1}$ and $E_{2}$, it is either the case that $E_{1}$ is before $E_{2}$, or simultaneous to $E_{2}$, or after $E_{2}$ [34]. Clearly, these relations are somehow ``````` `static”, i.e., they are not tensed and hold unchanged since ever and forever. For example, according to the B-theory, the extinction of the dinosaurs has always been, is and will always remain before the French revolution. Note however, that this relation seems to us to necessarily presume determinism, otherwise, in an indeterministic world, it would have not been true at a time earlier than the dinosaurs’ extinction itself actualized, because that extinction was not necessary in the first place (nor was the French revolution). At the remote past time in which both events ``````` `dinosaurs’ extinction” and ``````` `French revolution” were only potential events, their relation of before-after was also indeterminate. After the actualization of dinosaurs’ extinction but without an actualization of the French revolution, one rules out the possibility that the French revolution is before (or simultaneous to) the dinosaurs’ extinction, but it remains in general indeterminate whether the French revolution will ever actualize (although this may have a determined propensity [4]). Hence, B-theorists – who take the before-after relation as fixed – don’t derive determinism but assume it from the outset.
Regarding the metaphysical status of time, the B-theory naturally fits to the spacetime of (deterministic) classical theory and the theory of relativity, where events are organised on a 4-dimensional manifold. Moreover, B-theorists uphold Eternalism – i.e., the doctrine that all events are temporally on an equal footing and their relation of being in the past, present or future of each other – merely reflects their placement on the manifold [35]. Clearly, the concept of geometric time as expressed above falls into this B-theory camp.
On the contrary, A-theorists uphold the view that characterizes the positioning in time with a monadic attribution: given an event $E$, it is either the case that $E$ is present, or past, or future [34]. For A-theorists change really happens: ``````` `there is a way reality is (now, presently) which is complete [i.e., maximal given what exists (a pure state as physicist say)], but was different in the past and also will be different in the future”[35]. It is clearly this camp that to be more friendly to the concept of creative time explained above.
It seems that all A-theorists accept the theses of temporalism, i.e., that propositions change their truth-value over time, and of temporal disparity, i.e., the view that there is a metaphysical distinction between past, present and future [36]. The biggest disagreement is about which specific kind of temporal disparity one envisions. There are three main proposals: (i) Presentism [37, 38], which maintains that ``````` `no objects exist in time without being present” [35]; (ii) the Moving spotlight theory [39], an eternalist A-theory which takes events to be set on a fix manifold, but the present is singled out as if it was lit by a light moving along the manifold; and (iii) the Growing-block theory [40] (sometimes referred as the Growing-universe, or Growing-past), according to which events come into existence at present but, unlike in presentism, they persist in the past.
The moving spot-light, as an eternalist theory, does not seem compatible with creative time, showing that there is not a one-to-one correspondence between A-theories and the concept of creative time. However, temporalism is an essential ingredient for creative time to process and certain A-theories are the natural metaphysical framework for describing the processing of creative time. Let us focus on the the Growing-block view. More precisely, the latter states that the present is metaphysically privileged because it is then that events become determinate (or ``````` `real” in the standard parlance of philosophy of time, which we do not endorse). But as new events become present, the past ones remain equally ``````` `real”, resulting in a growing block of reality [35]. In particular, we find parallels with a particular version of Growing-block theory, namely, E. Barnes and R. Cameron’s growing cloud of determinacy theory [41], which upholds that the openness of the future has nothing to do with its alleged non-existence (a common conclusion in philosphy of time). In the words of the authors, ``````` `the future is as yet unsettled. [T]hink of this ‘unsettledness’ with respect to future states of the world as a type indeterminacy. For all times $t_{1}$ and $t_{2}$ such that $t_{2}$ is later than $t_{1}$, it is indeterminate at $t_{1}$ what the state of the world is at $t_{2}$[41].
This is the theory of the passage of time which seems most aligned with our concept of creative time. The future exists as a collection of real potentialities and the actualization of these potentialities is the act of creation that defines creative time. The present is therefore, for us, the edge of the events that turn from potential (the open future) to actual (the settled past). [^11] Although we have discussed the possibility that the remote past could again become indeterminate in Ref. [16]. So, while past, present and future are for us all in a sense real, the past is the (growing) collection of physical properties that got already actualized, the future is the collection of the potential properties, [^12] Note that some future properties have propensity one, hence are already actual [4]. and the present is the transition from potential to actual (see also [4]). [^13] Note that the present is not a global, universal property, but there are as many local presents as actualization events. Similar ideas were already put forward by H. Reichenbach who wrote: ``````` `The present, which separates the future from the past, is the moment when that which was undetermined becomes determined, and ’becoming’ means the same as ’becoming determined’.” [42]. Despite this, in Reichenbach, the present is still relative to someone or something, i.e. any moment can serve as the division between the determined past and the indeterminate future relative to who is simultaneous to that event (see [43] and reference therein for a discussion). Our proposal also has several similarities with S. McCall’s ``````` `objective time flow” [43]. Therein, the author considers that the universe has, at any point in time, a dynamical tree structure, with a single determined past and multiple possible futures. The concept of becoming is brought about by the update of the tree structure, when only one branch becomes determined. However, McCall’s proposal remains at an abstract level, whereas we present a naturalistic account of the indeterminate-actualization transition. Moreover, McCall does not conceive the possibility of objective (possibly biased) tendencies or propensities towards one possible future or another.
To conclude this section, we find it worth positioning our contribution within the metaphysical debate on time. First of all, although we advocate some form of Growing-block theory, we notice that most existing accounts thereof in the philosophy of time have been developed with an a priori metaphysical approach. The novelty of our approach is to provide a naturalistic characterization of the passage of time and of the present, as an objective transition from the past to the future. Moreover, one of the main criticisms against the standard Growing-block theory (the so-called skeptic objection), which rendered this approach quite unpopular among philosophers, is that since both past and present have the same ontological status (i.e., they exist, as opposed to the future), how can we know that we live in the present and not in the past? It seems that our ideas expressed above overcome this skeptical objection exactly due to its naturalistic character. In fact, although both past and present are determinate (and therefore different from the future), they are also different from each other. The present is when creative time processes, namely, when the determination (or actualization) happens and new information gets created. The present has a dynamical nature, whereas the past remains fixed.
VI More insights on time(s) from physical theories
VI.1 Relativity theory
As already recalled, geometric time is the conception of time derived from the regularities of some phenomena described by classical physics, like the motion of planets and satellites in the solar system which in turn led to a dominant deterministic world-view in physics. As remarked by G. E. M. Anscombe, ``````` `the high success of Newton’s astronomy was in one way an intellectual disaster: it produced an illusion… for this gave the impression that we had here an ideal of scientific explanation; whereas the truth was, it was mere obligingness on the part of the solar system, by having had so peaceful a history in recorded time, to provide such a model.” [44].
We agree with H. Reichenback that ``````` `the properties of time which the theory of relativity has discovered have nothing to do with its treatment as fourth dimension. This procedure was already possible in classical physics, where it was frequently used. However, according to the theory of relativity the four-dimensional manifold is of a new type; it obeys laws different from those of classical theory” [45]. And indeed it was the advent of the theory of (both special and general) relativity that popularized further the idea of geometric time. Therein, not only is time a parameter, but it seems to lose any special role with respect to space.
Time Physics Mathematics Logic Philosophy
Geometric Determinism Classical Non-tensed B-theory
Creative Indeterminism Intuitionistic Tensed A-theory
Table 1: Summary of the relation of concepts of creative and geometric time to differnt disciplines.
Geometric time intervals are not relativistically invariant. Indeed, the (geometric) time of a moving observer at speed $v$ (with respect to a given reference frame) undergoes a dilation of a factor $\sqrt{1-\frac{v^{2}}{c^{2}}}$, according to the Lorentz transformations. Geometric time can be manipulated by using the scenario of the twin paradox. This means to prepare two identical systems ($K$ and $L$), send one of them (say $L$) away, and finally bring $L$ back together with $K$. Since $L$’s local clock passes at a slower rate than the one of $K$, the geometric time elapsed for the system $L$ will be shorter than the one elapsed for $K$ (the higher $v$, the larger the dilation). Let us consider again the example of the weather in Geneva in a month time, and let us assume that the same computation of the outcome is carried out with a computer $K$ and with a moving computer $L$ (assuming that they are using the same algorithm and identical computers). If at the event in which they meet again, $K$’s clock has registered a time $\Delta t^{\prime}$ (which is the time that it takes for her computer to return an outcome), $L$’s program would still be running due to time dilation. This shows again that geometric time can be manipulated. On the contrary, it is unclear whether there is any way to let creative time ``````` `run faster”.
The question then naturally arises whether creative time, too, transforms relativistically. Looking at concrete experiments from particle physics, one would initially think that it necessarily does. Indeed, subatomic quantum particles like muons, undergo spontaneous decay which supposedly involves a genuine indeterministic process, and hence creative time. Muons have a mean lifetime of a only 2.2 µs, so on average they should be able to travel a half-survival distance of only about 456 meters without decaying [46]. The fact that we are able to detect cosmic rays muons (i.e., produced in the upper atmosphere) at sea level, however, means that they exist undecayed for a period (as seen from the reference frame of Earth) that is 25 times longer than their lifetime. Such an effect is explained by the fact that due to the muon’s velocity close to the speed of light, the muon lifetime undergoes relativistic time dilation. This could be considered a strong argument to show that also creative time transforms under the Lorentz transformations.
Yet, one can also think that the muon evolves deterministically while propagating through the atmosphere (thereby involving only geometric time) into a state of quantum superposition of decayed-undecayed and the process that actualizes the outcome ``````` `decayed” (which involves creative time) only happens inside the detector, hence at rest with Earth. In this way, creative time would not necessarily follow relativistic transformations. While this is admittedly a wild conjecture, it seems not to contradict the experimental evidence. This calls for further investigations.
One might think that geometric time is more friendly to relativity than creative time (for a discussion on this see also [15]). Indeed, if there were only geometric time, then everything would be given all at once (determinism) and there would be no difference between time and space. But, actually, creative time might also be quite friendly to (special) relativity. Consider a true random number generator that produces (creates) a bit and let us fix the coordinate system such that this bit is produced at the origin. Along a time-like line passing through the origin, there is a time when the bit’s value is indeterminate and a time when it is determinate. Next, consider a space-like line in the future. Assuming that the bit has a determined value only inside the future light cone (see [15]), one can “move” along that space-like line and notice, similarly to moving along the time-like line, a segment thereof where the bit’s value is indeterminate and a segment where it is determinate. Note that along the time-like line there are only 2 regions –one of indeterminacy and one of determinacy–, while along the space-like line there are 3 regions, i.e., one can “move” from indeterminate to determinate and (without “returning”) again to indeterminate.
Finally, the idea that creative time fundamentally separates the past from the future may prima facie seem add odds with relativity. However, the events of actualization that characterize the development of creative time are to be considered local therefore there is no (global) simultaneity of the present, in accordance with relativity. The event of actualization from indeterminate to determinate exists only within the forward light-cone, while remaining indeterminate in any other region of spacetime (see [15] for a thorough discussion). Therefore distant regions have different ticking of creative time. However, one can conjecture that due to quantum nonlocality (of the propensities that characterize the indeterminacy), different ``````` `local presents” may be synchronised by the actualization at a distance of a nonlocal propensity (see also [4]): yet another wild conjecture.
VI.2 Quantum theory
Quantum mechanics is the first physical theory in which indeterminism has been widely considered at a fundamental level. Indeed, Heisenberg uncertainty principle (which would be better named “indeterminacy”’ principle) and later on violation of Bell’s inequalities are strong indications that properties of quantum systems do not exist predetermined [47]. In this, quantum physics has inspired us to develop the concept of creative time, which happens at the moment when a quantum event, e.g. a measurement, determines a physical property (i.e. a single real outcome) that was previously only potential (as encapsulated by the probability amplitudes in the quantum state) [4, 48].
This can, however, also be applied in hindsight to classical theory that can consistently be interpreted as a fundamentally indeterministic theory [3, 10].
Creative time thus becomes associated with how, and under what circumstances the potentialities become actual, i.e. to the notorious measurement problem (also at the classical level, see [3, 4]). In fact, the concept of creative time is compatible only with these classes of interpretations that conceive a passage from the potential to the determinate [49]. In quantum theory these are the ``````` `spontaneous collapse theories” or interpretations , à la Copenhagen, where the collapse is induced by (some features of) the measurement. Other interpretations, such as the many-world, cannot accommodate the concept of creative time.
VI.3 Thermodynamics
A long standing problem that would be impossible not to mention is why do we ubiquitously observe a direction of time while the microscopic laws of physics are time-reversal invariant. This is commonly explained through statistical consideration, i.e., the fact that the thermodynamic quantity entropy cannot decrease in a closed system (second law of thermodynamics) [50]. This tension between the observed asymmetry in time and the time symmetry of the underlying microphysics are exemplified by the Loschmidt Paradox and by the Zermelo objection that Boltzmann’s H-Theorem is at odds with Poincaré recurrence theorem.
However, this is not necessarily the case even at the classical level, as discussed in detail in the works of B. Drossel and G. Ellis [8, 51, 52, 53, 54].
In a similar fashion, our distinction between creative and geometrical time can help clarify this matter. If there are fundamentally indeterminate events that require creative time, then there is also a fundamental asymmetry between past and future, as encapsulated by the second law. So, while admittedly this still requires to address the measurement problem (classical or quantum), paradoxes like the ones of Loschmidt or Zermelo are solved by merely noticing that Poincaré recurrence theorem simply fails [3].
VII conclusions
We join T. Maudlin when he states that ``````` `it is a fundamental, irreducible fact about the spatio-temporal structure of the world that time passes” [55]. However, the formalization of physical theories has to a large extent expelled the real passage of time from the description of natural phenomena (see also [1]). We have shown that to understand physics we should rethink our concept of time and distinguish between a geometric time (that appears in the deterministic equations of motion) and a creative time (that happens concurrently with events of actualization of potentialities).
In physics, indeed, we have related these two conceptions of time to, respectively, determinism and indeterminism. We then showed a parallel in classical and intuitionistic mathematics. The former takes a Platonistic approach by assuming that mathematical entities (such as real numbers) are given all at once, therefore avoiding the necessity of time. On the other hand, we showed that a naturalistic interpretation of intuitionistic mathematics requires a concept of creative time. Consequently, we have advocated a tensed multivalued logic (which can be true, false or indeterminate) for scientific propositions. Finally, we have related our discussion to known philosophical theories of time, showing in particular that geometric time fits best with the worldview of B-theory, whereas creative time with A-theory. Creative time has also led us to conceive a variation to the Growing-block theory, which however overcomes fundamental criticisms by providing a naturalistic account of the difference between past, present and future. These parallels in different disciplines are summarized in Table 1.
We conclude by stating that, in our view, physics is not only about sophisticated theories and fascinating technologies, it should also allow one to tell stories about “how nature does it”. But all stories require time [19]. Hence, we would like to rephrase a famous aphorism by French writer F. Rabelais as: “science without time is but ruin of intelligibility”. [^14] The original sentence reads: ``````` `Science without consciousness is but ruin of the soul”.
Acknowledgements.
We thank Barbara Drossel, George Ellis, Carl Posy, Cristian Mariani, Mauro Dorato, Tim Maudlin, Valerio Scarani, Clive Aw, Mark van Atten, Andrea Oldofredi, and Baptiste Le Bihan for useful discussions and for pointing out relevant literature. This research was supported by the FWF (Austrian Science Fund) through an Erwin Schrödinger Fellowship (Project J 4699), and the Swiss National Science Foundation via the NCCR-SwissMap.

### v2

@@
 Abstract
-We propose a distinction between two different concepts of time that play a role in physics: geometric time and creative time.
+We distinguish two different concepts of time that play a role in physics: geometric time and creative time.
 The former is the time of deterministic physics and merely parametrizes a given evolution.
 The latter is instead characterized by real change, i.e.
-novel information that gets created when a non-necessary event becomes determined in a fundamentally indeterministic physics.
-This allows us to give a naturalistic characterization of the present as the moment that separates the potential future from the determined past.
-We discuss how these two concepts find natural applications in classical and intuitionistic mathematics, respectively, and in classical and multivalued tensed logic, as well as how they relate to the well-known A- and B-theories in the philosophy of time.
+novel information that gets created when a non-necessary event becomes determinate in a fundamentally indeterministic physics.
+This allows one to give a naturalistic characterization of the present as the moment that separates the potential future from the determinate past.
+We discuss how these two concepts find natural applications in classical and intuitionistic mathematics, respectively, and in classical and intuitionistic logic, as well as how they relate to the well-known A- and B-theories in the philosophy of time.
+We acknowledge that we do not offer here a unified concept or a new philosophy of time.
+However, we contend that none of the existing philosophical accounts fully integrate both the geometric and creative concepts of time.
 I Introduction
-There is almost nothing that we perceive so ubiquitously than the passage of time.
+There is almost nothing that we perceive so ubiquitously as the occurrence of time.
 And yet our most successful physical theories still struggle to make sense of this concept in an unequivocal way.
@@
 However, in the words of I.
-Prigogine, ``````` `no formulation of the laws of nature that does not take into account this constructive role of time can ever be satisfactory” [2].
-To address this fundamental problem, we propose here two different concepts of time, which we call geometric time and creative time, respectively.
-We show that those stem from our fundamental assumption of physics as being either deterministic or indeterministic at the fundamental level, and that they both seem to contribute to our understanding of natural phenomena.
-In particular, we will see that geometric time is the parametric time appearing in the (deterministic) equations of motions of physics, just a coordinate that, together with the three spatial directions, labels fix events in a geometrical block-universe.
-On the other hand, creative time originates from the assumption that certain events which are fundamentally indeterminate become determinate.
-Our general approach is admittedly highly influenced by quantum mechanics, which has brought the concept of fundamental indeterminacy into the domain of physics.
-But we abstractly conceive the possibility that there is ontic indeterminacy in the world (also at the classical level, see [3, 4]).
-This idea of creative time resonates to some extent with the positions of the philosopher H.
-Bergson, well-known for having confronted Einstein about the nature of time, who stated: ``````` `Time is what prevents everything from being given all at once.
+Prigogine, ``no formulation of the laws of nature that does not take into account this constructive role of time can ever be satisfactory” [75].
+To address this fundamental problem, we distinguish here two different concepts of time, which we call geometric time and creative time, respectively.
+We show that those stem from our fundamental assumption of physics as being either deterministic or indeterministic at the fundamental, ontological level and that they both seem to contribute to our understanding of natural phenomena.
+In particular, we will see that geometric time is the parametric time appearing in the (deterministic) equations of motions of physics, just a coordinate that, together with the three spatial directions, labels fixed events in a geometrical block-universe.
+On the other hand, creative time arises from the assumption that certain events, which are fundamentally indeterminate, become determinate.
+Our general approach is admittedly influenced by quantum mechanics, which as a matter of fact brought the notions of fundamental indeterminacy and indeterminism into physics.
+More abstractly, we consider the possibility of ontic indeterminacy: a metaphysical feature of a physical quantity that exhibits “worldly imprecision” [62, 63, 4, 85].
+An ontologically indeterminate reading of the quantum state is, arguably, a widely held option (setting aside nonlocal hidden‐variable theories or many‐worlds).
+Although not a logical necessity, Heisenberg uncertainty together with violations of Bell inequalities provide strong reasons to regard quantum indeterminacy as more than a reasonable hypothesis.
+Moreover, we have shown in detail that, even at the classical level, if one adopts a finiteness‐of‐information principle, there is a consistent way to interpret classical theory as fundamentally indeterminate [46, 24, 29, 26].
+Throughout the paper, we therefore take seriously the working hypothesis of ontic (sometimes called fundamental) indeterminacy.
+In general, indeterminacy is a kinematic property of the state describing such a quantity.
+There is nothing subjectively uncertain here; in our naturalistic view, Nature simply has not yet fully determinate all objects.
+Indeterminism, on the other hand, refers to the possibility in Nature of having different potential (mutually exclusive) future states, given the present state and the laws of physics.
+Note that ontic indeterminacy in Nature is sufficient (though not necessary) for the existence of multiple potential future states, i.e., for indeterminism.
+Indeterminacy can be formalized in terms of the information existing about, for example, a certain physical property.
+In fact, we contend that the world is composed of material entities – e.g., particles and fields – but these do not exist simpliciter.
+What exists is structure: information about the arrangement of things, their relations, their positions, etc.
+This information is objective, existing in nature independently of any observer.
+It is physically instantiated in the degrees of freedom of material systems: a physicalist interpretation in the spirit of Landauer’s dictum that information is physical [52].
+More information can be created when a potential future is actualized, thereby reducing indeterminacy.
+It is the actualization of potentialities – when new information that previously did not exist is created – that makes creative time tick.
+This is in line with the ideas of mathematician A.
+Connes, who states: ”It’s quantum effervescence that generates the passage of time, not the other way around.” [22].
+This idea of creative time also resonates to some extent with the positions of the philosopher H.
+Bergson, well known for having confronted Einstein about the nature of time, who stated: ``Time is what prevents everything from being given all at once.
 […] It must therefore be development.
 Wouldn’t it then be the vehicle of creation and choice?
-Would time’s existence not prove that things are undetermined?
+Would time’s existence not prove that things are indeterminate?
 Would time not be this very indeterminacy?” [11].
-We somewhat twist this view around by assuming that it is indeterminacy that defines a different fundamental type of time.
-Indeed, if there is ontic indeterminacy in nature (i.e., physical pure states do not fully determine the future), this leads to multiple potential future states, i.e., to indeterminism and an open future.
-It is the actualization of these potentialities, i.e., when new information previously not existing is created, that makes this creative time tick.
+However, while vaguely inspired by Bergson, as physicists we twist the story around and adopt a more explicitly naturalistic view: it is the natural change in metaphysical indeterminacy (i.e., the creation of novel information and the passage from a potentiality to an actual property) that defines our creative time.
+We can thus say that, while geometric time “passes”, creative time “processes” or “occurs”, as the consequence of changes in the world.
+[^1] In the literature one often finds that time ``unfolds” or that time ``flows” to indicate that time is associated with something that really changes, and thus with creative time.
+Here we prefer to use the more active terms ``time occurs” or ``time processes”, because a film already shot does unfold, but nothing new is created there; whereas to flow is a somewhat misleading term, for it recalls the concept of speed, which already presupposes time.
 Figure 1: Typical evolution of a classical chaotic system in phase space.
-The initial state (a) has some ontic indeterminacy, i.e., the position and momentum are in this case not determined further than the finite region depicted in figure.
+The initial state (a) has some ontic indeterminacy, i.e., the position and momentum are in this case not determinate further than the finite region depicted in figure.
 When evolved through the deterministic equations of motions (b), the indeterminacy spreads (while preserving the same volume), leading to multiple potential future evolutions, i.e.
@@
 V).
-More specifically, the approach based on consecutive actualizations of potentialities, that is when creative time is at work, leads to a naturalistic characterization of the difference between past, present and future.
+More specifically, the approach based on consecutive actualizations of potentialities, that is when creative time is at work, leads to a naturalistic description of the difference between past, present, and future.
+We do not claim that the account we present here constitutes a new philosophy of time, nor that it resolves enduring debates, such as that between phenomenalism and scientism, regarding our perception or the intrinsic nature of time.
+Rather, we aim to emphasize the distinction between two different concepts of time, which we believe arise naturally from physical science and have parallels in various other disciplines.
 II Two times in physics
@@
 One can envision several procedures to reply to this question.
-On the one hand, it is possible to use the equations of motion of all the involved molecules of air, water vapour, etc., to calculate their evolution and so predict the weather in Geneva in a month time.
+On the one hand, it is possible to use the equations of motion of all the involved molecules of air, water vapor, etc., to calculate their evolution and so predict the weather in Geneva in a month’s time (thereby applying an algorithm, which is by definition a set of finite instructions).
 On the other hand, one can wait exactly one month, let the process develop, and finally observe the weather in Geneva.
-In-between, one could spend one week to collect more data – or merely waste a week – and only then start computing the evolution of all air molecules.
+In-between, one could spend one week to collect more data and only then start computing the evolution of all air molecules.
 Let us concentrate on the two first mentioned extremal procedures.
@@
 Before elaborating on that, we ought to distinguish between different fundamental interpretations of physics, which become especially relevant in dealing with complex systems like the one determining the weather in our example.
-The standard narrative (i) has it that classical physics is deterministic, i.e.
+The standard narrative, (i) has it that classical physics is deterministic, i.e.
 that each state is mapped one-to-one to all other possible states at all times (with time being the parameter that appears in the equations of motion).
-Although usually not stressed enough, this is based on two independent ingredients: the first is that since the equations of motion are ordinary differential equations, their solutions generally have a unique continuation in the parameter time, [^1] In exceptional cases, classical equations of motion of even simple systems do not have a unique (deterministic) solution, such as in the example of Norton’s dome [6].
-However, in highly complex systems – such as Lagrangian fluid particles in high Reynolds-number turbulence – the fact that deterministic trajectories are not unique is rather common (see [7] and references therein).
+We would like to stress that this is arguably based on a popular misconception, while, in fact, most physical theories can be regarded as indeterministic [36].
+In fact, the belief that classical physics is deterministic is based on two independent ingredients: the first is that since the equations of motion are ordinary differential equations, their solutions generally have a unique continuation in the parameter time, [^2] In exceptional cases, classical equations of motion of even simple systems do not have a unique (deterministic) solution, such as in the example of Norton’s dome [66].
+Moreover, in highly complex systems – such as Lagrangian fluid particles in high Reynolds-number turbulence – the fact that deterministic trajectories are not unique is rather common (see [41] and references therein).
 given appropriate initial conditions; the second is that the initial conditions exist fully determinate with infinite precision (i.e., a state is represented by a dimensionless mathematical point in phase space, in turn described by an n-tuple of real numbers).
@@
 In this case one can only statistically predict what are the possible future states compatible with the partial knowledge of the initial conditions.
-Finally, (iii) one can instead uphold the more radical assumption that the initial conditions (alternatively, the state of a system) simply do not exist determined with infinite precision, but there is an ontic indeterminacy in nature (for a collection of works on this see, [8, 3, 9, 10, 11, 4]; see also [12] and references therein for recent developments on metaphysical indeterminacy).
-In this case, although the equations of motion may remain unchanged, the system does not undergo a necessary evolution, but multiple future states are possible (perhaps with a determined objective tendency or propensity; see [4]), i.e., it is the ontic indeterminacy that leads to fundamental indeterminism (see Fig.
+Finally, (iii) one can instead uphold the more radical assumption that the initial conditions (alternatively, the state of a system) simply do not exist determinate with infinite precision, but there is an ontic indeterminacy in nature (for a collection of works on this see, [33, 24, 46, 48, 29, 26]; see also [20] and references therein for recent developments on metaphysical indeterminacy).
+In this case, although the equations of motion may remain unchanged, the system does not undergo a necessary evolution, but multiple future states are possible (perhaps with a determinate objective tendency or propensity; see [26]), i.e., it is the ontic indeterminacy that leads to fundamental indeterminism (see Fig.
 1).
@@
 The first procedure, under the ideal assumption of determinism (i), merely makes use of time as a parameter that simply indicates the coordinate of a certain evolution (e.g., sunny weather in one month from the start of the computation).
-Together with the three spatial coordinates (i.e., in Geneva in our example), this kind of time forms a manifold that contains all the events: a four-dimensional block universe picture where past, present and future are all on the same metaphysical footing (this is also known as Eternalism in the philosophy of time; see Sect.
+Together with the three spatial coordinates (i.e., in Geneva in our example), this kind of time forms a manifold that contains all the events: a four-dimensional block universe picture where past, present, and future are all on the same metaphysical footing (this is also known as Eternalism in the philosophy of time; see Sect.
 V).
 This can thus be called geometric time, the time that most physicists have come to think of as fundamental, both from classical physics and even more from relativity theory where time becomes “just” another dimension like space.
-In this sense, time is a relational label that allows one to position an event (with respect to another point of reference), but there is no real ``````` `becoming”.
+[^3] Though with a different signature in the metric.
+In this sense, time is a relational label that allows one to position an event (as earlier or later than to another point of reference), but there is no new information created in the world.
 On the other hand, let us consider the case of ontic indeterminacy in the physical state (iii).
-There, the actual development of a non-necessary physical process – such as the evolution of a highly chaotic system like the one that determines the weather, or of a quantum system – involves, in our opinion, a time that really ``````` `processes”.
-That is, a concept of time which sets an actual separation of the future from the past (with the present at their interface).
+There, the actual development of a non-necessary physical process – such as the evolution of a highly chaotic system like the one that determines the weather, or of a quantum system – involves, in our opinion, a time that really ``processes”.
+That is, a concept of time that sets an actual separation of the future from the past (with the present at their interface).
 And it does so when new information that was not existing before gets created – i.e., even having access to the whole existing information of the initial conditions (even of the whole state of the universe) before, would not determine all the future physical properties of a system afterwards.
 Again, we stress that we are not talking here about epistemic uncertainty, but rather we explicitly take a physicalist stance on information: it quantifies something objectively encoded in the degrees of freedom of the universe.
-In this sense information gets created when non-necessary physical events come about, independently of any observer.
-Note that, – and this is what happens also in quantum physics – although the state of the system is fully determined and it contains the maximal information about the systems (i.e., it is pure), this does not fully determine the future evolution (see also Ref.
+In this sense, information is created when non-necessary physical events obtain, independently of any observer.
+Note that – as in quantum physics – even given the pure state of a system – i.e., in Butterfield’s words, “a system’s maximal (or ‘complete’) set of intrinsic (or ‘possessed’) properties” [18] – this does not fully determine the future evolution (see also Ref.
 [62]).
-The concept of time that stems from this can thus be called creative time.
+Here, complete means that a state contains the maximal information presently available “in nature” about each potential property, i.e.
+the collection of all propensities.
+Yet only one potentiality, among its alternatives, eventually actualizes.
+The concept of time that arises from this can thus be called creative time.
 [^4] The distinction between geometric and creative time was sketched in Refs.
 [44, 25, 24, 81] and previously also in Ref.
-[17] where the creative time was called ``````` `historical time”.
+[49] where the creative time was called ``historical time”.
 On a similar note, G.
-Ellis remarked that ``````` `The most important property of time is that it unfolds.
-[…] The time that is the present at this instant will be the past at the next instant.
+Ellis remarked that ``The time that is the present at this instant will be the past at the next instant.
 […] It is this fundamental feature of time that is not encapsulated by today’s theoretical physics” [38] (see also [30]).
 Let us emphasize that the concept of creative time has nothing to do with language, nor with subjective or psychological processes; it ticks due to objective processes that happen by themselves in nature.
-Accordingly, there are two types of events: necessary events (e.g., when two geometric worldlines cross each other) and non-necessary events when potentialities actualize.
-Those instances of actualization determine the objective present.
-In other words, this fundamental “becoming” comes from the realization of one single outcome among the possible ones (with the contextual exclusion of the other previously valid futures) due to ontic indeterminacy.
-In the literature one often finds that time ``````` `unfolds” or that time ``````` `flows” to indicate that time is associated with something that really changes (so with creative time).
-Here we prefer to use the terms time ``````` `develops” or ``````` `processes”, because also a film already shot does unfold, but nothing new is created there; whereas to flow is a somewhat misleading term for it reminds of the concept of speed which has time already in it.
-So, henceforth, we will write that geometric time passes while creative time processes or happens.
-Coming back to the example of the weather in Geneva in a month, to carry out either of the envisioned procedures to determine the outcome takes some time.
-In the first procedure, i.e.
-when geometric time is at work, one maps the time that it takes to perform the computation $\Delta t^{\prime}$ to the time appearing in the equations of motion.
+Accordingly, there are two types of events: necessary events (e.g., when two geometric worldlines cross each other) and non-necessary events which occur when potentialities become actual.
+This is exemplified, for instance, by the case of a single photon impinging on an array of detectors—eventually, only one detector clicks.
+So, coming back to the example of the weather in Geneva in a month, one can carry out either of the envisioned procedures to determine the outcome – (i) compute the state of the weather in a month using the relevant laws of physics, or (ii) wait and directly observe the weather.
+Both these procedures take some time.
+In practice, to carry out the first procedure, one maps the time interval that it takes to perform the computation $\Delta t^{\prime}$ to the time appearing in the equations of motion.
 That is, to the interval $\Delta t=t_{f}-t_{i}$, where $t_{i}$ is the parametric time associated to the initial state and $t_{f}$ is the parametric time of the considered final state (i.e., the moment when we want to know the weather in Geneva).
-The (deterministic) computation lasts $\Delta t^{\prime}<\Delta t=1$ month, and returns as an outcome the weather in Geneva at $t_{f}$, say ``````` `sunny”.
+Note that both $\Delta t^{\prime}$ and $\Delta t$ are geometric times.
+For this to be a prediction, the (deterministic) computation should last $\Delta t^{\prime}<\Delta t=1$ month, and return as an outcome the weather in Geneva at $t_{f}$, say ``sunny”.
 [^5] In principle the time for the “prediction” could take longer than the time interval to observe the outcome.
 In such a case, one would anyways ask for consistency of the calculation afterwards.
-One can also think of more subtle questions about the weather, such as ``````` `will there be (anytime in history) 700 consecutive days of rain in Geneva?”.
+One can also think of more subtle questions about the weather, such as ``will there be (anytime in history) 700 consecutive days of rain in Geneva?”.
 If the complex system that forms the weather is fully deterministic, answering this question seems only a matter of geometric time, although the answer may remain undecided for any finite value of geometric time.
-One can run the computation of the dynamics of the system and at some point find that the answer is ``````` `yes”, if one finds the 700 consecutive days of rain, otherwise the algorithm will keep searching without ever halting (however, this question might be more complex, see Sect.
+One can run the computation of the dynamics of the system and at some point find that the answer is ``yes”, if one finds the 700 consecutive days of rain, otherwise the algorithm will keep searching without ever halting (however, this question might be more complex, see Sect.
 III).
-Two remarks are here in order.
-Firstly, note that $\Delta t^{\prime}$ can be speed up by a more efficient computation, namely, this depends on how fast one ``````` `runs” through the parameter time as it appears in the equations of motion (this depends on the used algorithm and, physically, one can really think of the speed of the head of a Turing machine in reading the tape).
-Moreover, time dilation in relativity can be seen as a physical process to modify the interval of geometric time (see Subsect.
+Note that $\Delta t^{\prime}$ can be sped up by a more efficient computation, namely, this depends on how fast one ``runs” through the parameter time as it appears in the equations of motion (this depends on the used algorithm and, physically, one can really think of the speed of the head of a Turing machine in reading the tape).
+Moreover, one can use any other effects that allow one to manipulate geometric time, such as time dilation in relativity that modifies the interval of geometric time $\Delta t^{\prime}$ (see Subsect.
 VI.1).
-So the map $\Delta t\rightarrow\Delta t^{\prime}$ is a one-to-many function.
-Secondly, once one reaches the final state with the associated outcome ``````` `sunny”, one gains full information about the initial state at $t_{i}$ and its consequences, but not necessarily the full information about the final state if there is some fundamental indeterminacy, i.e., the actual weather that will be observed may vary.
-On the contrary, the procedure involving creative time, namely, of waiting a month and carrying out an observation of the weather gives full information about the outcome, but provides in general little information about the initial state.
-This is because new information has been created along the way, through a series of indeterministic processes.
+So, the map $\Delta t\rightarrow\Delta t^{\prime}$ is a one-to-many function.
+In contrast, in the process involving creative time, then one has no choice but to wait a month and see what happens, because new information is created along the way through a series of indeterministic processes.
 So, the initial and the final states are not deterministically mapped to each other.
-The exactly same initial state under exactly the same physical conditions leads in general to different outcomes, so one can only infer a set of initial states that are compatible with the observed outcome.
-Creative time is thus associated to the possibility that physical properties are in general ontologically indeterminate (e.g., it is neither true that ``````` `the particle is within region $r$” nor that ``````` `the particle is not within $r$”, where $r$ is a region of space, see also Sect.
-IV).
 Creative time is required when an actualization event of the values of physical variables that were previously indeterminate occurs.
-It is the process of change from indeterminate to determinate – with an associated creation of novel information – that brings about this “happening” of time: creative time ticks when new information is created.
+It is the process of change from indeterminate to determinate – with an associated creation of novel information – that brings about this “occurrence” of time: creative time ticks when new information is created.
 III Two times in Mathematics
@@
 for each of these numbers there exists no algorithm – which is by definition a finite list of instructions – that outputs all their digits (they have infinite Kolmogorov complexity) [48].
-On the contrary, all the (irrational) numbers that we usually consider as prototypical examples of reals, such as $\sqrt{2}$ or $\pi$, are all computable (the ratio between the diagonal of a square to its edge is a simple algorithm that outputs all the digits of $\sqrt{2}$ and similarly the ratio of the circumference to the diameter of a circle for $\pi$).
-Obviously, all the rational numbers are also computable: they can be directly written down or compressed (in the case of repeating decimals) in a finite string.
-The standard way of regarding mathematics, the so-called classical-mathematics, is a form of Platonism which posits that, among the other mathematical entities, real numbers exist with their infinite series of digits, although there is in general no way to even label and thus grasp them (since there is no algorithm that generates them [20, 21].
-Hence, classical Platonistic mathematics is a timeless language, which per se is not a bug, but it becomes problematic when mathematics is elevated to the language of science, i.e., it is used to the describe the physical world.
-However, there are alternative approaches known as constructive mathematics, of which the most prominent is intuitionism.
-Therein, mathematical entities are not given at once, but rather are processes in development (digits get created continuously, one after the other).
-This provides mathematics with a concept of passage of time (and it would be only at infinite time – at the end of time, so to say – that the mathematical entities, such as real numbers, are completed into the ones defined by classical mathematics).
+On the contrary, the (irrational) numbers that we usually consider as prototypical examples of reals, such as $\sqrt{2}$ or $\pi$, are all computable (the ratio between the diagonal of a square to its edge is a simple algorithm that outputs all the digits of $\sqrt{2}$ and similarly the ratio of the circumference to the diameter of a circle for $\pi$).
+Obviously, all the rational numbers are also computable: they can be directly written down or compressed (e.g., in the case of repeating decimals) in a finite string.
+The standard way of regarding mathematics, the so-called classical-mathematics, is a form of Platonism which posits that, among the other mathematical entities, real numbers exist with their infinite series of digits, although there is in general no way to even label and thus grasp them (since there is no algorithm that generates them [45, 12]).
+Hence, classical Platonistic mathematics is a timeless language, which per se is not a bug, but it becomes problematic when mathematics is elevated to the language of science, i.e., is used to describe the physical world.
+However, there are alternative approaches known as constructive mathematics, of which one of the most prominent is intuitionism (see [15, 73, 10] and references therein).
+Therein, mathematical entities are not given at once, but rather are processes that develop in time (digits get created continuously, one after the other).
+[^6] More precisely, it is not necessarily new digits that come into existence, but the new information reduces the indeterminacy of intuitionistic real numbers [12].
+This explicitly provides mathematics with a concept of passage of time (and it would be only at infinite time – at the end of time, so to say [73] – that the mathematical entities, such as real numbers, are completed into the ones defined by classical mathematics).
+Indeed, one should stress that “the dependence of intuitionism on time is essential: statements can become provable in the course of time and therefore might become intuitionistically valid while not having been so before”.
+[50].
 The initiator of intuitionism, L.E.J.
-Brouwer, envisioned a ``````` `creative subject”, i.e., an idealized mind or mathematician, who is responsible for this progressive process of creation in time of, for instance, the digits of an uncomputable real numbers.
+Brouwer, envisioned intuitionistic mathematics as “having its origin in the perception of a move of time” [16].
+However, for him mathematics was the activity of the mind of a ``creative subject”, i.e., an idealized mind or mathematician, who is responsible for this progressive process of creation in time of, for instance, the digits of an uncomputable real numbers [16].
+Note that, for Brouwer, time is merely internal, i.e., it is not physical time at all – neither geometric nor creative in the sense discussed here (see [9]).
 Several authors, however, have distanced themselves from this controversial concept of a creative subject.
-In particular, one of us (N.G.) has put forward the idea that the digits of a typical real number are generated by a true random number generator, [^4] More precisely, it is not necessarily new digits that come into existence, but the new information reduces the indeterminacy of intuitionistic real numbers [21].
-i.e., a natural process that is able to create a piece of new information by changing a fundamentally indeterminate bit into a determined one [10].
+In particular, one of us (N.G.) has put forward the idea that the digits of a typical real number are generated by a “natural” random number generator (see also footnote 6), i.e., a natural process – as opposed to the action of an agent – that creates new information in the world by changing a fundamentally indeterminate bit into a determinate one [47].
 [^7] For more works that relate constructive mathematics to physics, see [45] and [42].
-This version of intuitionism could thus be labeled ``````` `objective” or ``````` `naturalistic” intuitionism.
-In this way, typical real numbers become a graspable concept directly linked to creative time, that is, to the change from the indeterminate to the determinate.
+This version of intuitionism, which departs from the standard view (see e.g., [73, 86]), could thus be labeled ``objective” or ``naturalistic” intuitionism, as characterized in [8].
+In this way, typical real numbers become a graspable concept directly linked to creative time, that is, to the change from the indeterminate to the determinate happening in natural phenomena.
 Hence, naturalistic intuitionism, is a tensed mathematical language.
@@
 In fact, it is possible to think that the full information about those numbers is contained in the (finite) algorithm that defines them, i.e., in their initial conditions.
-For instance, one can ask what is the 43800th digit of $\pi$ and the answer is given by running an algorithm that outputs the digits of $\pi$ and picking its 43800th digit.
-Note that it is not necessary to go through all the previous digits of $\pi$ to compute 43800th one.
+For instance, one can ask what is the 43800th decimal place of $\pi$ and the answer is given by running an algorithm that outputs the digits of $\pi$ and picking its 43800th decimal.
+The answer happens to be a 9, and it is easily computable because it is not necessary to go through all the previous digits of $\pi$ to compute 43800th one.
 There are so-called digit-extraction algorithms that allow one to directly compute the nth digit of $\pi$ [70].
@@
 Digit-extraction algorithms, also in base 10, are known for several other irrational computable real numbers, such as for $e$.
-This can be seen as a further evidence that computable (irrational) numbers are already fully determined and do not require creative time.
-The point is that since this is found through a deterministic outcome, the answer to this kind of questions is fully contained in the algorithm, which can be run more or less fast, giving the ability to manipulate this ``````` `mathematical geometric time”.
+This can be seen as a further evidence that computable (though irrational) numbers are already fully determinate and do not require creative time.
+The point is that since this is found through a deterministic outcome, the answer to this kind of question is fully contained in the algorithm, which can be run more or less fast, giving the ability to manipulate this ``mathematical geometric time”.
 This exactly resembles the weather example in Sect.
@@
 There are no algorithms compressing the information of that number.
-The only existing ``````` `algorithm” is the number itself.
-Each next digit is generated by a genuinely natural random process, therefore, asking which is the 43800th digit of such a number requires to wait for all 43800 instances of creating of the ``````` `next digit” (see footnote 4), i.e., it requires ``````` `mathematical creative time”.
-Finally, one can also ask a question of the kind ``````` `are there 700 consecutive sevens in the digits of a number.
-[^7] The example of 700 consecutive sevens is borrowed from C.
+The only existing ``algorithm” is the number itself (more precisely there is no finite list of operations that generates the number, but the only way to indicate the number is to list its infinite digits).
+Each next digit is generated by a genuinely random natural process.
+Therefore, asking what the 43800th digit of such a number requires waiting through all 43800 instances of the creation of the ``next digit” (see footnote 6); that is, it requires ``mathematical creative time”.
+In intuitionism, these numbers are generated in time by what are usually called non-lawlike choice sequences, i.e., sequences of computable numbers with new bits (or digits) created as time processes.
+[86, 73, 10, 64, 65].
+[^9] In what follows, we will consider that uncomputable numbers are generated by non-lawlike sequences, in contrast to computable numbers that are generated by lawlike sequences.
+In our naturalistic version of intuitionism [8], fresh bits get generated by a natural random number generator, i.e., a process occurring in nature that is able to produce true randomness, and a choice sequence $\alpha(n)$ is the outcome of a computable function that takes as an argument $\alpha(n-1)$ – where $n$ is the time index – and all the random bits previously generated (assumed to be independent and identically distributed).
+[^10] Note that although the computable function deterministically produces outcomes given suitable inputs, the inputs themselves are random—thus rendering the output indeterminate.
+Typical, uncomputable real numbers can be regarded as generated through this natural process of creation of new information as creative time processes.
+This has a connection to the mathematical model of indeterminate physics that was introduced in Ref.
+[24].
+There, the standard real number are substituted by so-called finite information quantities (FIQs), which only encode the propensity of the digits of a physical quantity to actualize.
+FIQs can be regarded as a particular species of choice sequences, where the next generated bit is biased by taking the majority vote of the values taken by the previous $k$ bits, with $k$ being an odd integer.
+In this framework, the propensities emerging in FIQs stem from the inherent randomness of $50\%-50\%$ bit choices within these sequences.
+If a strong majority of recent $k$ bits is 1, it is likely that this trend will persist in the near future.
+This likelihood is captured as a propensity, constrained to rational values.
+(see [47] for a detailed discussion).
+[^11] To fully recover the comparison between physics and mathematical intuitionism, one might expect a discussion extending beyond the notion of choice sequences.
+The present work, however, focuses on a more foundational stage: establishing the meaningfulness of formulating physical theory using choice sequences, particularly in light of a constructive reconsideration of time.
+Finally, one can also ask a question of the kind ``are there 700 consecutive sevens in the digits of a number.
+[^12] This example is borrowed from C.
 Posy [73].
@@
 $\pi$, the question is interesting.
-The situation is similar to the example of the weather – i.e., 700 consecutive days of rain if one assumes determinism – but the axiomatic character of mathematics makes the solution to this question clearer.
-One can, in fact, think of programming two softwares: The first, $S_{1}$ is exactly the one we used in the physical example of the weather, i.e., the systematic search of a 700 consecutive sevens in $\pi$; on the other hand, the second software, $S_{2}$ outputs all the theorems that can be derived within Peano arithmetic and it halts should it find the negation of the conjecture.
-If the statement is true, $S_{1}$ will find this sequence of 700 sevens and therefore halt after a finite (geometric) time.
-If the statement is provable (within Peano’s arithmetic), $S_{2}$ will halt after a finite (though presumably enormous) geometric time.
-Interestingly, if one assumes that neither halts, then the statement is false (if not $S1$ would halt), but not provable within Peano’s arithmetic (if not $S_{2}$ would halt).
-[^8] For the sake of completeness, there is also the fourth logical possibility that both $S_{1}$ and $S_{2}$ halt, a case which would prove Peano’s arithmetic inconsistent.
+The situation is similar to the example of the weather – i.e., whether there will be 700 consecutive days of rain if one assumes determinism.
+One can, in fact, think of programming two types of software: The first, $S_{1}$, is exactly the one we used in the physical example of the weather, i.e., the systematic search of 700 consecutive sevens in $\pi$; on the other hand, the second software, $S_{2}$ outputs all the theorems that can be derived within Peano arithmetic and it halts should it find the negation of the conjecture that there are 700 consecutive sevens in $\pi$.
+If the conjecture is true, $S_{1}$ will find such a sequence of 700 sevens and therefore halt after a finite (geometric) time.
+If the conjecture is provably false (within Peano’s arithmetic), $S_{2}$ will halt.
+Interestingly, if neither halts, then the statement is false (otherwise, $S1$ would halt), but this is unprovable within Peano’s arithmetic (otherwise, $S_{2}$ would halt).
+[^13] Given that the lifetime of the universe is sufficient for carrying out the calculation.
+Moreover, for the sake of completeness, there is also the fourth logical possibility that both $S_{1}$ and $S_{2}$ halt, a case which would prove Peano’s arithmetic inconsistent.
 This would be an example of Gödel’s celebrated theorem.
-Note that mathematicians don’t use softwares like $S_{1}$ and $S_{2}$, but use their creativity to find shortcuts to analyse such statements, i.e., de facto they use creative time.
+Note that mathematicians don’t use software like $S_{1}$ and $S_{2}$, but use their creativity to find shortcuts to analyse such statements, i.e., de facto they use creative time.
 In the case of the weather in Geneva, if one assumes that the system is indeterministic, like uncomputable numbers, then there is no choice but wait and see.
-However, if one trusts the equations describing the weather evolution and the fully determined initial conditions, then one could run analogues of $S_{1}$ and $S_{2}$, as for computable numbers.
+However, if one trusts the equations describing the weather evolution and the fully determinate initial conditions, then one could run analogues of $S_{1}$ and $S_{2}$, as for computable numbers.
 Note however, that this assumes one can compress the initial conditions into a finite algorithm (hence into a computer which is necessarily finite).
+We would like to emphasize that one of the key features of intuitionism is that, while introducing time and finiteness into mathematics, it still preserves the concept of the continuum.
+This makes intuitionism a suitable mathematical framework for physics, as most physical theories rely on continuity (e.g., Lorentz transformations in relativity require the continuum).
+However, the intuitionistic continuum is “sticky” or “viscous” [73] because, at any finite time, choice sequences are still ongoing, and one cannot select a specific point from the continuum, as it is not yet determinate whether the sequence will converge to above or below or on the specific point.
+This is the origin of Brouwer’s famous continuity theorem that states that all total functions , i.e., functions defined for every element of its domain, are continuous [47].
 Let us acknowledge that this distinction between geometric and creative time in mathematics distances our view from the most accepted position in intuitionism, such as that of C.
@@
 van Atten [87, 88], who consider all numbers, whether computable or not, as generated by new information that comes about, so that they all require creative time in our parlance.
+Exactly how these approaches to intuitionism diverge, philosophically, is a question we have only recently begun to investigate with B.
+Bentzen [8] and that will require more systematic work.
 IV Logic
-[…]
-Figure 2: Representation of ontic indeterminacy in phase space.
-Now, the principle of non-contradiction imposes already that propositions about physics must be referred to a time.
-Indeed, if one states simpliciter that ``````` `the system is inside $r$” and ``````` `the system is not inside $r$”, this is a violation of the principle of non-contradiction, since we do observe in nature the same system being sometimes inside $r$ and sometimes not.
-One has then to refer to a time to qualify the propositions.
-If one adds ``````` `at time $t$” to both propositions, then the principle of non-contradiction holds because –at any time $t$ – only one of these propositions would be true and the other not.
-On the contrary, if one labels one proposition with ``````` `at time $t_{1}$” and the other with ``````` `at time $t_{2}$”, with $t_{1}\neq t_{2}$, it is possible that they are both true.
-Hence, the principle of non-contradiction applies only to propositions at the same time.
-Note that this time is not necessarily creative, but it could just be the label appearing in the equations of motion, i.e., geometric time.
-Coming instead to a scenario where there is ontic indeterminacy, we now show that the law of the excluded middle fails.
-Consider for instance a model of physics (classical, quantum, or even post-quantum) in which a (pure) state is represented by a finite volume, $R$, in phase-space (as opposed to a dimensionless mathematical point).
-That is, there is not merely a lack of knowledge about the system (as it is the case in, e.g., statistical physics), but this state contains the complete information existing at present about all the properties of the system (see also [3, 4]).
-Since, as we have already argued, this ontic indeterminacy generally leads to indeterminism in the future evolution of a system, [^10] Except for integrable classical dynamical systems.
-statements about future properties are in general neither true nor false because they are yet to be determined.
-And it is this process of determination that requires creative time.
-In general, in the case of ontic indeterminacy, one needs to introduce a multivalued tensed logic according to which a proposition is either true or false only at a time $\tau\geq\tau_{0}$ when it has become determined, but its truth value was ``````` `indeterminate” (as a third logical value) before $\tau_{0}$ (for multivalued logic see [28, 29]).
-As an example, consider a true random number generator that outputs a (single determined) bit of information, say $b=0$.
-That event defines the (creative) time instant $\tau_{0}$.
-After such an event that brings into existence this new bit of information, the truth value of the proposition ``````` `b=0” is true and the proposition ``````` `b=1” is false, but before $\tau_{0}$ both propositions had a truth value ``````` `indeterminate”.
-This represents a breakdown of the principle of the excluded middle.
-Note that such a failure of the excluded middle is found in all constructive mathematics – e.g.
-intuitionism [25].
-However, despite these many similarities, we emphasize that our approach differs substantially from ``````` `orthodox” intuitionism.
-First of all, as already remarked in Section III, we reject the Brouwerian notion of ``````` `creative subject” in intuitionistic logic, advocating for a naturalistic version of intuitionism in which new information is created in the universe by natural processes.
-Moreover, contrarily to what we have advocated, we stress that orthodox intuitionistic logic is not three-valued (nor multivalued for that matter) – as proven by Gödel [30]– and that logical propositions are typically timeless.
-Finally, consider a system whose present state is represented by a finite volume in phase space $R$ (see Fig.
-2).
-This means that although the state is fully determined (but not as a collection of real numbers [9]), it does not fully determine the future.
-In this case, a statement like ``````` `the system is located in phase space within $r\subset R$” may lead to think that the principle of the excluded middle fails also for statements about the present.
-However, such a statement is not indeterminate, but actually false, for it is simply not the case that the system is located inside $r$ in phase space.
-The only presently true statement about the state of the system is ``````` `the system is located at $R$ in phase space” and nothing more.
-On the contrary, the statement ``````` `the system will be (at some future time) located in phase space at $r$” is currently indeterminate, because it is not impossible that the system will become more localized (in phase space).
-Once the evolution of the system develops, the statement will become either true or false.
-This process is what requires creative time.
-Hence, the failure of the excluded middle concerns only future statements, where the future is identified by a change of creative time.
+We have seen that creative time emerges, both in physics and in mathematics, from indeterminacy, i.e., from the possibility of having new information that comes into existence.
+Another way to investigate the role of time and indeterminacy is to consider propositions about physical properties and analyze their logical structure.
+In this logical framework, one formulates physical properties as propositions; for example, $p(R)$ denotes the proposition “the system has property RR”.
+The aim of this approach is to assign truth values to physical propositions and to combine them using logical connectives.
+In particular, $p(R)$ is true, at the empirical level, if and only if an ideal (dichotomic) measurement of the corresponding property of the system would yield the outcome $R$ with certainty.
+[^14] By “dichotomic” we mean that the measurement can only answer, ideally, whether the system has property $R$ or not.
+In a classical physical theory, this criterion can be implemented by identifying $R$ with a region of phase space and checking whether the system’s state is entirely contained in $R$.
+Two central principles uphold in classical logic are the law of non-contradiction, $\neg(p\wedge\neg p)$, and the law of the excluded middle (LEM), $p\vee\neg p$.
+We now examine how different notions of time influence the logical structure of propositions about physical properties, and how these two fundamental principles are affected in relation to underlying ontologies, whether deterministic or indeterministic.
+Determinism in logic can be defined as follows: “If $p$ is true then it has always been the case that $p$ will be true”, for all propositions $p$.
+Indeterminism is just its negation: $\exists p$ such that $p$ is true and it has not always been the case that $p$ will be true” [84].
+In what follows, we discuss two main arguments that, in our opinion, hold in logic when applied to physics: (i) the law of non-contradiction requires a (geometric) time label for both determinism and indeterminism; and (ii) the law of the excluded middle fits best with deterministic logic (and hence with geometric time), whereas indeterminism (related to creative time) fits best with the rejection of the LEM.
+(i) Law of non-contradiction
+Since physical states evolve in time, propositions about physical properties, such as $p(R)$ above, must reflect this temporality.
+In fact, if one asserts simpliciter – without any temporal qualification – that “the system has property RR” and also that “the system does not have property RR”, one is led to an apparent contradiction.
+For a system whose state changes in time, each statement may be correct at different instants; hence, without a time index, both $p(R)$ and $\neg p(R)$ would be true, thereby violating the principle of non-contradiction.
+A natural remedy is to supply a temporal index.
+Once one writes “the system has property $R$ at time tt” and “the system does not have property $R$ at time tt”, the principle of non-contradiction is respected, since at any given time only one of these propositions can be true (the pure state of a system is unique at any given time).
+By contrast, if the propositions concern different instants, $t_{1}\neq t_{2}$, both may be true without inconsistency.
+Thus the principle of non-contradiction applies only to propositions referring to the same time, a point already emphasized by Kant [51].
+Hence, time is necessary for logical consistency when dealing with propositions referring to physical systems.
+In deterministic physics, this time can simply be the parameter appearing in the equations of motion, i.e., geometric time.
+In indeterministic physics, by contrast, one must consider the logical structure of propositions in which creative time plays a role.
+(ii) Law of the excluded middle
+The law of the excluded middle (LEM) sits naturally within a logically deterministic framework, in which the truth values of all propositions are fixed and constant in time.
+By contrast, in a fundamentally indeterministic physics—where future contingencies are not yet determinate and become so only at a later moment of creative time—logic models in which the LEM may fail are more naturally motivated, as already noted by Aristotle (see [92], Ch.
+9).
+In such a setting, the truth values of propositions about the future need not be determinate in the present.
+To capture this situation, one may adopt non-classical logical frameworks that explicitly accommodate indeterminacy.
+In the remainder of this section, we discuss some of these approaches, which illustrate the failure of the LEM and provide formal frameworks aligned with indeterminacy and the associated notion of creative time.
+We stress that we do not intend here either to give a systematic review of these suitable logical systems, nor to advocate for any specific one.
+Our main aim is to give notable examples from the literature that align with our goal of introducing temporality into logic in a way that the LEM does not hold, for the description of indeterministic physics.
+We will now show that this can be achieved in different ways, such as by introducing temporality explicitly into the syntax (like in Prior’s tense logic where a future operator is introduced as a new logical connective), or solely into the semantics (like in Kripke models in intuitionistic logic).
+A first possibility for dealing with indeterminate propositions would be to adopt a multivalued logic – with more truth values than just true and false, for example an “indeterminate” value – as in the influential work of J.
+Łukasiewicz [55].
+This approach, however, raises other foundational difficulties, such as the controversial definition of negation, and will therefore not be discussed further here.
+On the other hand, tense logic – when combined with an open-future and certain semantics – offers a way to maintain a two-valued logic while allowing a failure of the LEM for future contingents.
+No third truth value is introduced: propositions are always either true or false relative to a given time and a given future branch.
+However, in an open-future setting, propositions about the future are evaluated with respect to multiple admissible branches.
+As a result, a future-tense proposition may be true on some branches and false on others, and thus the LEM need not hold globally for future contingents, even though bivalence is preserved locally on each branch.
+[^15] Strictly speaking, tense logic does not reject the LEM at the syntactic level; the failure arises at the semantic level under branching-time (open-future) interpretations, which we assume throughout.
+This perspective was first formalized by A.
+Prior (see [67] and references therein) within a modal framework introducing temporal operators such as the future operator $F$, where $Fp$ is read as “the proposition $p$ will be true at some future time relative to the present.” Note that although $F$ introduces temporality explicitely in the syntax, it is not necessarily the case that this leads to a violation of LEM unless one assumes an open-future semantics of a certain kind [^16] In fact, Prior initially explored systems that allow violations of the law of non-contradiction while retaining the LEM [71], but he later developed the Peircean and Ockhamist systems [76, 67], in which non-contradiction is preserved and the LEM fails specifically for future contingents under branching-time semantics.
+On this note, an important development of Prior’s ideas is the so-called branching-time semantics [7, 67].
+In a branching model of time, truth at the present moment is evaluated relative to a set of admissible future branches.
+At the present time ($t=0$), FpFp is true if and only if, for every admissible future branch $b$, there exists a future time $t_{b}>0$ along that same branch at which $p$ holds.
+Conversely, $\neg Fp$ is true if and only if, for every admissible future branch $b$ and for every future time $t>0$ along that branch, $p$ fails to hold.
+When $p$ holds along some branches but not others, neither $Fp$ nor $\neg Fp$ is true at present, and the law of excluded middle $Fp\lor\neg Fp$ fails for such future contingents, even though no third truth value is introduced.
+Here we do not commit to any particular logical system; rather, we generically advocate for a two-valued logic in which the law of excluded middle fails for future contingencies, while the principle of non-contradiction is preserved.
+We take this approach because, in our view, it best captures the notion of creative time in an indeterministic physics (again, this is similar to Aristotle’s position [92], however motivated in naturalistic terms).
+Note that such a failure of the law of excluded middle also appears in constructive mathematics—e.g., intuitionism [73, 50]—making this proposal simultaneously meet our desiderata for a logic model and remain consistent with our perspective on mathematics in Sect.
+III.
+In fact, Kripke models for intuitionistic logic can be understood temporally if we view possible worlds as construction stages that track the mental activity of Brouwer’s constructing subject [91], or as time steps associated with actualization in our naturalistic view [8].
+Note, however, that temporality is not made explicit in the syntax, as it is in tense logic [91].
+Let us analyze in more detail the failure of the LEM in this view.
+In our indeterministic framework, physical properties will eventually actualize, yielding eventual bivalence: either $p$ or $\neg p$ will be true once the relevant (creative) time has occurred.
+However, this eventual bivalence does not imply that either $p$ or $\neg p$ is already true at the present time.
+In intuitionistic logic, to say that $p$ is true at $t=0$ means that $p$ can be proven at $t=0$, using the information available at that time.
+in the context of indeterministic physics, “can be proven” is understood as “is in fact already determinate”, as argued in greater detail in [8].
+Accordingly, for the disjunction appearing in the LEM to be true, the proposition $p\vee\neg p$ must satisfy two conditions: (i) either $p$ or $\neg p$ is true, and (ii) it is determinate which of the two alternatives holds.
+[^17] In standard (i.e., non-naturalistic) intuitionism, this condition is formulated epistemically: $p\vee\neg p$ is true iff (i) either $p$ or $\neg p$ is true, and (ii) it is known, or has been proven, which alternative holds.
+Since we assume that propositions about the future are, in general, ontically indeterminate, the LEM does not hold for them at the present time.
+It is precisely the process of determination – which restores the LEM for present propositions – that requires creative time.
+It may also be useful to consider intuitionistic free logic, since it allows to deal with objects that may not yet exist, such as worlds or branches that did not yet actualized [72].
+We emphasize again that our approach differs from “orthodox” intuitionism, for we reject the Brouwerian notion of the “creative subject” in intuitionistic logic, advocating instead a naturalistic version of intuitionism in which new information is created in the universe by natural processes [8].
+[^18] There is historical evidence that Brouwer discussed the possibility of sequences generated by a physical process, e.g., throwing a die [89, 90].
+Although this is a matter of historical debate, it seems that Brouwer rejected this possibility, “probably because it was based on the physical world” [16].
+In summary, even in logic we see natural connections between different logic models and determinism versus indeterminism.
+We have argued that classical logic is most compatible with a deterministic worldview and aligns naturally with geometric time, whereas tense logic and (naturalistic) intuitionistic logic—where the LEM fails—provide a better fit for indeterminism and the notion of creative time.
 V Relation to the philosophy of time
-Before discussing in detail how our proposal of creative and geometric times relates to some positions in the philosophy of time, we would like to stress that, as most physicists, we are not taking a Humean position.
-On the contrary, we believe that laws of nature are real causal connections (they have an ontological status) and that they guide the evolution of things of the world at different times.
-Note that these laws may not connects states at different times by necessity (determinism), but only state a propensity [4].
-In either case, this puts an emphasis on the role of time in physical laws.
-In the vast literature devoted to the philosophy of time (see, e.g., [31, 32] and references therein), one finds a main separation into two camps, following the seminal works of J.
+In the vast literature devoted to the philosophy of time (see, e.g., [93, 74, 19, 53, 54] and references therein), one finds a main separation into two camps, following the seminal works of J.
 M.
@@
 The two camps are customarily referred to as A-theory and B-theory of time.
-While we acknowledge that the distinction in A- and B-theories presents a variety of modification and it may not even encapsulate the most striking features of time in physics (e.g., directionality, openness of the future, objective becoming, etc.), we cannot avoid drawing some parallel with our distinction between creative and geometric time, respectively.
+While we acknowledge that the distinction in A- and B-theories presents a variety of modifications and it may not even encapsulate the most striking features of time in physics (e.g., directionality, openness of the future, objective becoming, etc.), we cannot avoid drawing some parallel with our distinction between creative and geometric time, respectively.
+[^19] Admittedly, there are a large variety of A- and B-theories.
+Here we believe we follow some widely accepted understandings thereof.
 In fact, starting in reversed order, B-theorists regard positioning in time in dyadic relational terms: given two events $E_{1}$ and $E_{2}$, it is either the case that $E_{1}$ is before $E_{2}$, or simultaneous to $E_{2}$, or after $E_{2}$ [68].
-Clearly, these relations are somehow ``````` `static”, i.e., they are not tensed and hold unchanged since ever and forever.
-For example, according to the B-theory, the extinction of the dinosaurs has always been, is and will always remain before the French revolution.
-Note however, that this relation seems to us to necessarily presume determinism, otherwise, in an indeterministic world, it would have not been true at a time earlier than the dinosaurs’ extinction itself actualized, because that extinction was not necessary in the first place (nor was the French revolution).
-At the remote past time in which both events ``````` `dinosaurs’ extinction” and ``````` `French revolution” were only potential events, their relation of before-after was also indeterminate.
-After the actualization of dinosaurs’ extinction but without an actualization of the French revolution, one rules out the possibility that the French revolution is before (or simultaneous to) the dinosaurs’ extinction, but it remains in general indeterminate whether the French revolution will ever actualize (although this may have a determined propensity [4]).
-Hence, B-theorists – who take the before-after relation as fixed – don’t derive determinism but assume it from the outset.
-Regarding the metaphysical status of time, the B-theory naturally fits to the spacetime of (deterministic) classical theory and the theory of relativity, where events are organised on a 4-dimensional manifold.
-Moreover, B-theorists uphold Eternalism – i.e., the doctrine that all events are temporally on an equal footing and their relation of being in the past, present or future of each other – merely reflects their placement on the manifold [35].
-Clearly, the concept of geometric time as expressed above falls into this B-theory camp.
-On the contrary, A-theorists uphold the view that characterizes the positioning in time with a monadic attribution: given an event $E$, it is either the case that $E$ is present, or past, or future [34].
-For A-theorists change really happens: ``````` `there is a way reality is (now, presently) which is complete [i.e., maximal given what exists (a pure state as physicist say)], but was different in the past and also will be different in the future”[35].
-It is clearly this camp that to be more friendly to the concept of creative time explained above.
-It seems that all A-theorists accept the theses of temporalism, i.e., that propositions change their truth-value over time, and of temporal disparity, i.e., the view that there is a metaphysical distinction between past, present and future [36].
-The biggest disagreement is about which specific kind of temporal disparity one envisions.
-There are three main proposals: (i) Presentism [37, 38], which maintains that ``````` `no objects exist in time without being present” [35]; (ii) the Moving spotlight theory [39], an eternalist A-theory which takes events to be set on a fix manifold, but the present is singled out as if it was lit by a light moving along the manifold; and (iii) the Growing-block theory [40] (sometimes referred as the Growing-universe, or Growing-past), according to which events come into existence at present but, unlike in presentism, they persist in the past.
+Clearly, these relations are somehow ``static”, i.e., they are not tensed and hold unchanged since ever and forever.
+On the contrary, A-theorists uphold the view that characterizes the positioning of events in time with a monadic attribution: given an event $E$, it is either the case that $E$ is present, or past, or future [68].
+Concerning the metaphysical status of time, the B-theory aligns more naturally with the spacetime structure of (deterministic) classical physics and relativity theory, where events are organized on a 4-dimensional manifold.
+Moreover, B-theorists tend to endorse Eternalism, i.e., the doctrine that all events are temporally on an equal footing and their relation of being in the past, present or future of each other merely reflects their placement on the manifold [40] (but, as we will see, this is not necessarily the case).
+Clearly, the concept of geometric time as expressed above fits best with the worldview offered by the B-theory camp.
+A first point about Eternalism is that they seem to us to tacitly assume that events exist predeterminate, since all events necessarily are – in the static sense of being part of the block universe.
+As an example, consider two historically well-known events: the extinction of the dinosaurs and the French Revolution.
+In an indeterministic world, neither of these events was necessary; therefore, their before-after relation, as conceived in B-theories, would supposedly relate events that might never occur.
+At a remote past time, when both events – “dinosaurs’ extinction” and “French Revolution” – were merely potential, their before–after relation was likewise indeterminate, simply because the events themselves were not bound to happen.
+After the actualization of the dinosaurs’ extinction, but before the actualization of the French Revolution, one can rule out the possibility that the French Revolution is before (or simultaneous with) the extinction of the dinosaurs.
+However, it remains generally indeterminate whether the French Revolution will ever actualize (even if this indeterminacy may be characterized by a definite propensity [26]).
+[^20] We are not considering here the further complication that, from an evolutionary perspective, if the French Revolution occurs, this entails that humans have evolved, and thus the event must necessarily follow the extinction of the dinosaurs.
+Hence, B-theorists, who take the before–after relation as fixed, seem to us to assume that events are bound to occur.
+Moreover, we note that the idea of ontic indeterminacy, which gives rise to an open future, is in tension with B-theories taken at face value.
+In fact, B-theorists deny the existence of the metaphysical properties of pastness, presentness, and futurity.
+For instance, in a greatly influential work to support a B-theory view in the philosophy of time, D.
+H.
+Mellor states that ``nothing really is past, present or future in itself”, and that ``tense is not an aspect of reality” [61].
+In particular, we stress that B-theorists reject the idea that the future exists in itself.
+And by denying the existence of the future altogether, they also cannot accommodate the idea of an open future: There is no such a thing as the ``future” and, ipso facto, no ``open future”.
+Hence, geometric time – which, as we have argued, is closely tied to determinism in physics – finds a more natural alignment with B-theory.
+In this view, there is no ontological distinction between past, present, and future, and nothing genuinely new ever comes into existence: no new information is added to the world.
+A-theorists, on the contrary, accept the thesis of temporalism, i.e., that propositions change their truth-value over time, and of temporal disparity, i.e., the view that there is a metaphysical distinction between past, present and future [23].
+This leads to the view that changes really happen: ``there is a way reality is (now, presently) which is complete [i.e., maximal given what exists (a pure state as physicist say)], but was different in the past and also will be different in the future”[40].
+A-theory allows for a genuinely open future and fundamental indeterminism, which makes it the framework most naturally aligned with the notion of creative time.
+The biggest disagreement among different A-theories is about which specific kind of temporal disparity one envisions.
+There are three main proposals: (i) Presentism [57, 77, 94], which maintains that ``no objects exist in time without being present” [40]; (ii) the Moving spotlight theory [21], an eternalist A-theory which takes events to be set on a fix manifold, but the present is singled out as if it was lit by a light moving along the manifold; and (iii) the Growing-block theory [43] (sometimes referred as the Growing-universe, or Growing-past), according to which events come into existence at present but, unlike in presentism, they persist in the past.
 The moving spot-light, as an eternalist theory, does not seem compatible with creative time, showing that there is not a one-to-one correspondence between A-theories and the concept of creative time.
-However, temporalism is an essential ingredient for creative time to process and certain A-theories are the natural metaphysical framework for describing the processing of creative time.
-Let us focus on the the Growing-block view.
-More precisely, the latter states that the present is metaphysically privileged because it is then that events become determinate (or ``````` `real” in the standard parlance of philosophy of time, which we do not endorse).
-But as new events become present, the past ones remain equally ``````` `real”, resulting in a growing block of reality [35].
-In particular, we find parallels with a particular version of Growing-block theory, namely, E.
+The view we endorse here is akin in spirit to the Growing-block view, though with important distinctions.
+The growing block states that the present is metaphysically privileged because it is then that events become determinate (or ``real” in the standard parlance of philosophy of time, which we do not endorse).
+But as new events become present, the past ones remain equally ``real”, resulting in a growing block of reality [40].
+In particular, we find inspiration in a variant of the Growing-block theory, namely, E.
 Barnes and R.
-Cameron’s growing cloud of determinacy theory [41], which upholds that the openness of the future has nothing to do with its alleged non-existence (a common conclusion in philosphy of time).
-In the words of the authors, ``````` `the future is as yet unsettled.
-[T]hink of this ‘unsettledness’ with respect to future states of the world as a type indeterminacy.
+Cameron’s growing cloud of determinacy theory [6], which upholds that the openness of the future has nothing to do with its alleged non-existence (a common conclusion in philosophy of time).
+In the words of the authors, ``the future is as yet unsettled.
+[T]hink of this ‘unsettledness’ with respect to future states of the world as a type of indeterminacy.
 For all times $t_{1}$ and $t_{2}$ such that $t_{2}$ is later than $t_{1}$, it is indeterminate at $t_{1}$ what the state of the world is at $t_{2}$[6].
-This is the theory of the passage of time which seems most aligned with our concept of creative time.
-The future exists as a collection of real potentialities and the actualization of these potentialities is the act of creation that defines creative time.
-The present is therefore, for us, the edge of the events that turn from potential (the open future) to actual (the settled past).
+As such, this theory aligns well with our physics-inspired idea that the future exists, metaphysically different from the past, as a collection of real potentialities and the actualization of these potentialities is the act of creation that defines creative time.
+However, the Growing-block picture, along with the growing cloud of determinacy theory, is tenable only if we assume the existence of an absolute global present – a statement that holds in classical physics but is rejected by the theory of relativity.
+Therefore, our version of a growing-block universe incorporates the necessary modifications to remain compatible with relativistic principles.
+The present is defined, for us, as the edge of the events that turn from potential (the open future) to actual (the settled past).
 [^21] Although we have discussed the possibility that the remote past could again become indeterminate in Ref.
 [81].
-So, while past, present and future are for us all in a sense real, the past is the (growing) collection of physical properties that got already actualized, the future is the collection of the potential properties, [^12] Note that some future properties have propensity one, hence are already actual [4].
+For our view to remain compatible with relativity (see Section VI.1 and Ref.
+[25]), we must relax the absoluteness of determinate facts: the actualization of a propensity should itself be a relative property, defined with respect to regions of space–time, yet without invoking observers.
+Hence, a bit that becomes determinate does so only within its future light cone, while it remains indeterminate elsewhere.
+This results in multiple local presents, each defined by the collection of determination events lying on the edges of their past light cone.
+The set of local presents, however, contrarily to the Growing-block picture, cannot be collected into a global present.
+We note that these considerations resonate with an approach to quantum gravity due to R.
+Sorkin, named causal sets [83, 32, 31].
+This is an approach that takes spacetime as fundamentally made of discrete events related by partial order relations which encapsulate causality.
+The dynamics is a ``growth process” in which new elements come into existence from their causal ancestors.
+Yet, our view allows to maintain that past, present and future are all in a sense real, metaphysically distinct, but relative to a light cone: the past is the (growing) collection of physical properties that got already actualized, the future is the collection of the potential properties, [^22] Note that some future properties have propensity one, hence are already actual [26].
 and the present is the transition from potential to actual (see also [26]).
-[^13] Note that the present is not a global, universal property, but there are as many local presents as actualization events.
 Similar ideas were already put forward by H.
-Reichenbach who wrote: ``````` `The present, which separates the future from the past, is the moment when that which was undetermined becomes determined, and ’becoming’ means the same as ’becoming determined’.” [42].
-Despite this, in Reichenbach, the present is still relative to someone or something, i.e.
-any moment can serve as the division between the determined past and the indeterminate future relative to who is simultaneous to that event (see [43] and reference therein for a discussion).
+Reichenbach who wrote: ``The present, which separates the future from the past, is the moment when that which was indeterminate becomes determinate, and ’becoming’ means the same as ’becoming determinate’.” [78].
+In Reichenbach, however, the present is still relative to an observer or to a reference frame; that is, any moment can serve as the division between the determinate past and the indeterminate future, relative to whoever is simultaneous with that event (see [59] and references therein for a discussion).
 Our proposal also has several similarities with S.
-McCall’s ``````` `objective time flow” [43].
-Therein, the author considers that the universe has, at any point in time, a dynamical tree structure, with a single determined past and multiple possible futures.
-The concept of becoming is brought about by the update of the tree structure, when only one branch becomes determined.
-However, McCall’s proposal remains at an abstract level, whereas we present a naturalistic account of the indeterminate-actualization transition.
+McCall’s ``objective time flow” [59].
+Therein, the author considers that the universe has, at any point in time, a dynamical tree structure, with a single determinate past and multiple possible futures.
+The concept of becoming is brought about by the update of the tree structure, when only one branch becomes determinate.
+However, McCall’s proposal remains at an abstract level, whereas we present a naturalistic (physical) account of the indeterminate-actualization transition.
 Moreover, McCall does not conceive the possibility of objective (possibly biased) tendencies or propensities towards one possible future or another.
-To conclude this section, we find it worth positioning our contribution within the metaphysical debate on time.
-First of all, although we advocate some form of Growing-block theory, we notice that most existing accounts thereof in the philosophy of time have been developed with an a priori metaphysical approach.
-The novelty of our approach is to provide a naturalistic characterization of the passage of time and of the present, as an objective transition from the past to the future.
-Moreover, one of the main criticisms against the standard Growing-block theory (the so-called skeptic objection), which rendered this approach quite unpopular among philosophers, is that since both past and present have the same ontological status (i.e., they exist, as opposed to the future), how can we know that we live in the present and not in the past?
-It seems that our ideas expressed above overcome this skeptical objection exactly due to its naturalistic character.
+We find it worth positioning our contribution within the metaphysical debate on time.
+First of all, the novelty of our approach is to provide a naturalistic characterization of the passage of time and of the present, as an objective transition from the past to the future.
+Moreover, one of the main criticisms against the Growing-block theory (the so-called skeptic objection), which rendered this approach quite unpopular among philosophers, is that since both past and present have the same ontological status (i.e., they exist, as opposed to the future), how can we know that we live in the present and not in the past?
+It seems that our ideas expressed above overcome this skeptical objection exactly due to its naturalistic character, based on indeterministic physics.
 In fact, although both past and present are determinate (and therefore different from the future), they are also different from each other.
@@
 The present has a dynamical nature, whereas the past remains fixed.
+To conclude, we have shown that geometric time is more naturally aligned with the worldview of B-theories, in which all events exist as well-determinate within a block universe.
+By contrast, creative time fits best with A-theories, which allow for an open future and thus fundamental indeterminism.
 VI More insights on time(s) from physical theories
@@
 M.
-Anscombe, ``````` `the high success of Newton’s astronomy was in one way an intellectual disaster: it produced an illusion… for this gave the impression that we had here an ideal of scientific explanation; whereas the truth was, it was mere obligingness on the part of the solar system, by having had so peaceful a history in recorded time, to provide such a model.” [44].
+Anscombe, ``the high success of Newton’s astronomy was in one way an intellectual disaster: it produced an illusion… for this gave the impression that we had here an ideal of scientific explanation; whereas the truth was, it was mere obligingness on the part of the solar system, by having had so peaceful a history in recorded time, to provide such a model.” [2].
 We agree with H.
-Reichenback that ``````` `the properties of time which the theory of relativity has discovered have nothing to do with its treatment as fourth dimension.
+Reichenback that ``the properties of time which the theory of relativity has discovered have nothing to do with its treatment as fourth dimension.
 This procedure was already possible in classical physics, where it was frequently used.
@@
 And indeed it was the advent of the theory of (both special and general) relativity that popularized further the idea of geometric time.
-Therein, not only is time a parameter, but it seems to lose any special role with respect to space.
+Therein, not only is time a parameter, but it seems to lose any special conceptual role with respect to space (besides the signature of the Minkowski metric, which assigns opposite signs to spatial and temporal components).
 Time Physics Mathematics Logic Philosophy
-Geometric Determinism Classical Non-tensed B-theory
-Creative Indeterminism Intuitionistic Tensed A-theory
+Geometric Determinism Classical With LEM B-theory
+Creative Indeterminism Intuitionistic Without LEM A-theory
 Table 1: Summary of the relation of concepts of creative and geometric time to differnt disciplines.
@@
 Since $L$’s local clock passes at a slower rate than the one of $K$, the geometric time elapsed for the system $L$ will be shorter than the one elapsed for $K$ (the higher $v$, the larger the dilation).
-Let us consider again the example of the weather in Geneva in a month time, and let us assume that the same computation of the outcome is carried out with a computer $K$ and with a moving computer $L$ (assuming that they are using the same algorithm and identical computers).
+Let us consider again the example of the weather in Geneva in a month’s time, and let us assume that the same computation of the outcome is carried out with a computer $K$ and with a moving computer $L$ (assuming that they are using the same algorithm and identical computers).
 If at the event in which they meet again, $K$’s clock has registered a time $\Delta t^{\prime}$ (which is the time that it takes for her computer to return an outcome), $L$’s program would still be running due to time dilation.
 This shows again that geometric time can be manipulated.
-On the contrary, it is unclear whether there is any way to let creative time ``````` `run faster”.
+On the contrary, it is unclear whether there is any way to let creative time ``run slower”.
 The question then naturally arises whether creative time, too, transforms relativistically.
@@
 This could be considered a strong argument to show that also creative time transforms under the Lorentz transformations.
-Yet, one can also think that the muon evolves deterministically while propagating through the atmosphere (thereby involving only geometric time) into a state of quantum superposition of decayed-undecayed and the process that actualizes the outcome ``````` `decayed” (which involves creative time) only happens inside the detector, hence at rest with Earth.
+Yet, one can also think that the muon evolves deterministically while propagating through the atmosphere (thereby involving only geometric time) into a state of quantum superposition of decayed-undecayed and the process that actualizes the outcome ``decayed” (which involves creative time) only happens inside the detector, hence at rest with Earth, similarly to an atom in a superposition of an excited and a ground state.
+If the atom ``decays” (gets de-excited) then a photon is produced and the detection of the latter can demonstrate that the measurement took place.
+But if a photon is not detected, the atom can still be in superposition until it interacts with a detector.
 In this way, creative time would not necessarily follow relativistic transformations.
@@
 This calls for further investigations.
-One might think that geometric time is more friendly to relativity than creative time (for a discussion on this see also [15]).
-Indeed, if there were only geometric time, then everything would be given all at once (determinism) and there would be no difference between time and space.
+One might think that geometric time is more friendly to relativity than creative time (for a discussion and proposed counterarguments see also [13, 25]).
 But, actually, creative time might also be quite friendly to (special) relativity.
-Consider a true random number generator that produces (creates) a bit and let us fix the coordinate system such that this bit is produced at the origin.
+Consider a true random number generator that produces (creates) a bit and let us fix the coordinate system such that this bit is produced at a certain spacetime location, say the origin for simplicity.
 Along a time-like line passing through the origin, there is a time when the bit’s value is indeterminate and a time when it is determinate.
-Next, consider a space-like line in the future.
-Assuming that the bit has a determined value only inside the future light cone (see [15]), one can “move” along that space-like line and notice, similarly to moving along the time-like line, a segment thereof where the bit’s value is indeterminate and a segment where it is determinate.
+Next, consider a space-like line that intersects the future light cone (see Fig.
+2).
+Assuming that the bit has a determinate value only inside the future light cone (see [25]), one can “move” along that space-like line and notice, similarly to moving along the time-like line, a segment thereof where the bit’s value is indeterminate and a segment where it is determinate.
 Note that along the time-like line there are only 2 regions –one of indeterminacy and one of determinacy –, while along the space-like line there are 3 regions, i.e., one can “move” from indeterminate to determinate and (without “returning”) again to indeterminate.
-Finally, the idea that creative time fundamentally separates the past from the future may prima facie seem add odds with relativity.
-However, the events of actualization that characterize the development of creative time are to be considered local therefore there is no (global) simultaneity of the present, in accordance with relativity.
-The event of actualization from indeterminate to determinate exists only within the forward light-cone, while remaining indeterminate in any other region of spacetime (see [15] for a thorough discussion).
-Therefore distant regions have different ticking of creative time.
-However, one can conjecture that due to quantum nonlocality (of the propensities that characterize the indeterminacy), different ``````` `local presents” may be synchronised by the actualization at a distance of a nonlocal propensity (see also [4]): yet another wild conjecture.
+Figure 2: Illustration of determinacy propagation within the light cone. Dotted lines indicate states of indeterminacy, while solid lines represent determinacy.
+Finally, the idea that creative time fundamentally separates the past from the future may prima facie seem at odds with relativity.
+However, the events of actualization that characterize the occurrence of creative time are to be considered local therefore there is no (global) simultaneity of the present, in accordance with relativity.
+In Ref.
+[25], we have flashed out a proposal to make relativity and fundamental indeterminism compatible.
+This is to assume that event of actualization from indeterminate to determinate exists only within the forward light-cone, while remaining indeterminate in any other region of spacetime [25].
+(In)determinacy is, therefore, itself relative, i.e., determinate facts are not absolute.
+This idea has gained considerable relevance in recent analyses in the foundations of quantum theory [14, 17, 28].
+Our proposal, however, develops a theory of relative (determinate) facts with respect to regions of spacetime rather than to observers, as is the case in most of the above references.
+To understand this, consider for example two distant, space-like separated, observers, Alice and Bob, each with a local natural random number generator.
+If Alice’s random number generator actualizes a previously indeterminate bit to, say, the value $a=0$.
+For Bob, however, the value of the bit $a$ remains indeterminate until the intersection of the future light cones of Alice and Bob.
+In this light, the metaphysical property of being determinate is objective, but relative to a certain region of spacetime.
+Therefore, different regions have different ticking of creative time.
+However, one can conjecture that due to quantum nonlocality (of the propensities that characterize the indeterminacy), different ``local presents” may be synchronized by the actualization at a distance of a nonlocal propensity (see also [26]).
+This admittedly remains a serious challenge and likely calls for a fully relativistic treatment of quantum nonlocality.
 VI.2 Quantum theory
-Quantum mechanics is the first physical theory in which indeterminism has been widely considered at a fundamental level.
-Indeed, Heisenberg uncertainty principle (which would be better named “indeterminacy”’ principle) and later on violation of Bell’s inequalities are strong indications that properties of quantum systems do not exist predetermined [47].
+Quantum mechanics is the first physical theory in which indeterminism has been widely – though not universally – regarded as fundamental.
+Indeed, Heisenberg uncertainty principle (which would be better named “indeterminacy”’ principle) and, later on, the violation of Bell’s inequalities provided strong indications that properties of quantum systems do not exist predeterminate [69].
 In this, quantum physics has inspired us to develop the concept of creative time, which happens at the moment when a quantum event, e.g.
@@
 In fact, the concept of creative time is compatible only with these classes of interpretations that conceive a passage from the potential to the determinate [56].
-In quantum theory these are the ``````` `spontaneous collapse theories” or interpretations , à la Copenhagen, where the collapse is induced by (some features of) the measurement.
-Other interpretations, such as the many-world, cannot accommodate the concept of creative time.
+In quantum theory these are the ``spontaneous collapse theories” or interpretations, à la Copenhagen, where the collapse is induced by (some features of) the measurement.
+Other interpretations – such as the many-world and Bohmian mechanics – cannot accommodate the concept of creative time.
+It should be noted that, while indeterminacy in classical and quantum theory leads to some common features, it also presents fundamental differences (see Ref.
+[27] for a detailed discussion).
+In quantum mechanics, certain physical quantities are incompatible (represented by noncommuting operators), and their indeterminacies are linked through Heisenberg’s uncertainty relations—for example, position and momentum.
+In contrast, while classical physics can be supplemented with noncontextual hidden variables (such as the real numbers used in the theory), any completion of quantum physics would require contextual and nonlocal additional variables.
+In philosophical terms, classical indeterminacy is considered shallow, whereas quantum indeterminacy is deep [63, 27].
 VI.3 Thermodynamics
@@
 This tension between the observed asymmetry in time and the time symmetry of the underlying microphysics are exemplified by the Loschmidt Paradox and by the Zermelo objection that Boltzmann’s H-Theorem is at odds with Poincaré recurrence theorem.
-However, this is not necessarily the case even at the classical level, as discussed in detail in the works of B. Drossel and G. Ellis [8, 51, 52, 53, 54].
+However, this is not necessarily the case even at the classical level, as discussed in detail in the works of B.
+Drossel and G.
+Ellis [33, 34, 35, 37, 39].
 In a similar fashion, our distinction between creative and geometrical time can help clarify this matter.
@@
 So, while admittedly this still requires to address the measurement problem (classical or quantum), paradoxes like the ones of Loschmidt or Zermelo are solved by merely noticing that Poincaré recurrence theorem simply fails [24].
+The consequences of indeterminacy – and the role of creative time – in statistical mechanics, arguably a domain where we possess a physical explanation of time’s direction, require further analysis and are the subject of ongoing work.
 VII conclusions
 We join T.
-Maudlin when he states that ``````` `it is a fundamental, irreducible fact about the spatio-temporal structure of the world that time passes” [55].
+Maudlin when he states that ``it is a fundamental, irreducible fact about the spatio-temporal structure of the world that time passes” [58], or, in our jargon, that time occurs.
 However, the formalization of physical theories has to a large extent expelled the real passage of time from the description of natural phenomena (see also [82]).
-We have shown that to understand physics we should rethink our concept of time and distinguish between a geometric time (that appears in the deterministic equations of motion) and a creative time (that happens concurrently with events of actualization of potentialities).
-In physics, indeed, we have related these two conceptions of time to, respectively, determinism and indeterminism.
+We have shown that to understand physics in indeterministic terms – and we argued that there is strong conceptual and empirical ground to do so – we should rethink our concept of time and distinguish between a geometric time (that appears in the deterministic equations of motion) and a creative time (that processes concurrently with the events of actualization of potentialities).
+In physics, indeed, we have tried to relate these two conceptions of time to, respectively, determinism and indeterminism.
 We then showed a parallel in classical and intuitionistic mathematics.
-The former takes a Platonistic approach by assuming that mathematical entities (such as real numbers) are given all at once, therefore avoiding the necessity of time.
-On the other hand, we showed that a naturalistic interpretation of intuitionistic mathematics requires a concept of creative time.
-Consequently, we have advocated a tensed multivalued logic (which can be true, false or indeterminate) for scientific propositions.
-Finally, we have related our discussion to known philosophical theories of time, showing in particular that geometric time fits best with the worldview of B-theory, whereas creative time with A-theory.
-Creative time has also led us to conceive a variation to the Growing-block theory, which however overcomes fundamental criticisms by providing a naturalistic account of the difference between past, present and future.
+The former takes a Platonistic approach by assuming that mathematical entities (such as real numbers) are given all at once, therefore avoiding the necessity of creative time.
+On the other hand, we showed that a naturalistic interpretation of intuitionistic mathematics requires a concept of creative time [8].We have therefore advocated a set of desiderata for a logical framework that seem to best capture indeterminism and creative time: namely, a logic that incorporates temporality (either syntactically or purely semantically) and that, by adopting an open-future semantics, rejects the law of the excluded middle.
+Finally, we have related our discussion to some philosophical theories of time, showing in particular that geometric time fits best with the worldview of B-theories, whereas creative time with A-theories.
+Creative time has also led us to conceive a variation to the Growing-block theory, which however overcomes fundamental criticisms by providing a naturalistic account of the difference between past, present, and future.
+Notably, we showed that to be consistent with relativity, there are many local presents that cannot be collected into a global present.
 These parallels in different disciplines are summarized in Table 1.
@@
 Rabelais as: “science without time is but ruin of intelligibility”.
-[^14] The original sentence reads: ``````` `Science without consciousness is but ruin of the soul”.
+[^23] The original sentence reads: ``Science without consciousness is but ruin of the soul”.
 Acknowledgements.
-We thank Barbara Drossel, George Ellis, Carl Posy, Cristian Mariani, Mauro Dorato, Tim Maudlin, Valerio Scarani, Clive Aw, Mark van Atten, Andrea Oldofredi, and Baptiste Le Bihan for useful discussions and for pointing out relevant literature.
+We thank Clive Aw, Bruno Bentzen, Jeremy Butterfield, Mauro Dorato, Barbara Drossel, George Ellis, Michael Esfeld, Baptiste Le Bihan, Cristian Mariani, Tim Maudlin, Andrea Oldofredi, Carl Posy, Valerio Scarani, Christian Wüthrich and Mark van Atten, for useful discussions and for pointing out relevant literature.
 This research was supported by the FWF (Austrian Science Fund) through an Erwin Schrödinger Fellowship (Project J 4699), and the Swiss National Science Foundation via the NCCR-SwissMap.